### v2

# DeCoF: Generated Video Detection via Frame Consistency

Abstract
The escalating quality of video generated by advanced video generation methods leads to new security challenges in society, which makes generated video detection an urgent research priority. To foster collaborative research in this area, we construct the first open-source dataset explicitly for generated video detection, providing a valuable resource for the community to benchmark and improve detection methodologies. Through a series of carefully designed probe experiments, our study explores the significance of temporal and spatial artifacts in developing general and robust detectors for generated video. Based on the principle of video frame consistency, we introduce a simple yet effective detection model (DeCoF) that eliminates the impact of spatial artifacts during generalizing feature learning. Our extensive experiments demonstrate the efficacy of DeCoF in detecting videos produced by unseen video generation models and confirm its powerful generalization capabilities across several commercial proprietary models.
1 Introduction
Following the success of diffusion models in generating highly realistic images (Nichol & Dhariwal, 2021; Rombach et al., 2022; Parmar et al., 2023), a growing number of researchers have shifted their focus to the advancement of these models for video generation (Kawar et al., 2023; Blattmann et al., 2023; Zhang et al., 2023). Such developments have turned diffusion models into robust and versatile video generators, enabling the creation of high-quality videos from general conditional inputs, such as text and images. While this powerful and flexible generation capacity has the potential to alleviate the stress of daily life and inspire creativity in work, it concurrently introduces a new crisis of trust (Sharma et al., 2023; Barrett et al., 2023). Therefore, there is an urgent need for detectors capable of identifying generated videos.
In this paper, we propose a challenging task: Generated Video Detection, which aims to create universal detectors to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models. To evaluate generated video detectors, we create the first comprehensive generated video detection dataset: the $GeneratedVideoForensics$ (GVF) dataset. The GVF dataset encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments. Moreover, pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos. Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures (Khachatryan et al., 2023; Zhang et al., 2023; Wang et al., 2023a; Sterling., June 2023), such as pixel-space or latent-space models, which aims to measure the generality of the generated video detectors.
Figure 1: Can the detector trained on a specific video generation model cope with the challenges presented by unseen generation models?
Considering that video content contains individual frame contents and smooth consistency across multiple frames, a natural idea is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) models to achieve generated video detection. Our exploratory experiments found that these models have excellent discriminative capabilities for specific T2V models but cannot be applied to detect unseen T2V models.
Refer to caption (a) Geometry
Refer to caption (b) Layout
Refer to caption (c) Frequency
Refer to caption (d) Color
(e) Temporal
Figure 2:  Spatial and temporal artifacts on generated video. We illustrate the spatial and temporal artifacts present in the generated videos. Spatial artifacts: (a) errors in geometric appearance, (b) errors in image layout, (c) frequency inconsistency, (d) color mismatch; Temporal artifacts: (e) mismatch between frames.
Further probe experiments in Table 1 reveal that the spatiotemporal convolutional models predominantly identify video authenticity by detecting spatial artifacts, resulting in generated video detection that has degenerated into a 2D problem. A critical question arises: Spatial and temporal artifacts, which artifacts are key to constructing universal detectors?
To explore this problem, we first analyze the possibility of building a universal generated video detector through spatial artifacts. We convert the 3D problem of generated video detection into a 2D problem, generated image detection, and implement T2V detection based on existing text-to-image (T2I) detection methods. In section  5.2, we evaluate the performance of these generated image detectors on generated video detection, noting that the reason why these detectors lack generality to unseen generative models is that capturing specific spatial artifacts to identify video authenticity, while spatial artifacts are specific to the T2V model, these detectors are unable to capture more general and high-level spatial artifacts to establish better decision boundaries. So far, we conclude that spatiotemporal convolutional models primarily identify video authenticity by detecting spatial artifacts but are unable to detect more general spatial artifacts. Since spatial artifacts are easily detected, learning of temporal artifacts is difficult to perform.
To rectify this biased learning approach, mitigate the adverse effects induced by spatial artifacts, and construct a more generalizable generated video detector by capturing temporal artifacts, we propose to decouple temporal and spatial artifacts by mapping video frames to feature space, where the distance between features is negatively correlated with image similarity, through a pre-trained CLIP:ViT (Dosovitskiy et al., 2020). More importantly, the features of real video frames are still continuous after being mapped by the mapping function with the above requirements. Meanwhile, the frames of the generated video with inter-frame mismatch will have anomalies in their features. Besides, it greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned. Specifically, we only implement the learning of inter-feature anomalies through a temporal artifact modeling module with two transformer layers (Vaswani et al., 2017). Extensive exploring experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection, especially significant generalization capabilities for unseen generation models.
Our contributions are summarized as follows:
  • We present the first dataset that is comprehensive and scalable as new models emerge for benchmarking generated video detectors.
  • We explore the possibility of detecting generated videos by capturing spatial and temporal artifacts and point out the drawbacks of the detector by capturing spatial artifacts and their causes.
  • We develop a universal generated video detector named DeCof, validating through extensive experiments, including tests on commercial proprietary models.
2 Related work
2.1 Generation Models
Recently, with the success of diffusion models in generating images (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021; Esser et al., 2021; Rombach et al., 2022; Peebles & Xie, 2023), a number of Research explores text-to-video Diffusion Models (VDMs). Some work (Ramesh et al., 2022; Blattmann et al., 2023; Khachatryan et al., 2023; Wang et al., 2023a)ground their models on latent-based VDMs, for example, Text2Video-zero (Khachatryan et al., 2023) generates videos by using motion dynamics to enhance the latent code of a text-to-image stable diffusion model without requiring training on any video data, and ModelScopeT2V (Wang et al., 2023a) extends it with T2I stable diffusion model of convolutional and attentional blocks and trained using image-text and video-text datasets, ZeroScope (Sterling., June 2023) is a watermark-free T2V generative model trained with the weights of ModelScopeT2V, which is optimized for 16:9 compositions. On the other hand, PYoCo (Ge et al., 2023)and Make-A-Video (Singer et al., 2022) anchor their models on pixel-based VDMs. Show1 (Zhang et al., 2023) integrates pixel-based and latent-based VDMs, using Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, while latent-based VDMs upscale these low-resolution videos from pixel-based VDMs to then create high-resolution videos with low computation cost.
2.2 Generated Content Detections
In recent years, a large amount of work has focused on generated image detection (McCloskey & Albright, 2019; Zhang et al., 2019; Barni et al., 2020; Corvi et al., 2023; Sha et al., 2023), especially for unseen generation models. A simple classifier trained by Wang et al. (Wang et al., 2020) on pagan-generated images using multiple data augmentations can be generalized to 20 unseen other GAN-generated images. Frank et al. (Frank et al., 2020) concluded that the artifacts of the generated images are caused by the upsampling method. Wang et al. (Wang et al., 2023b) identify generated images through the difference of images reconstructed by a pre-trained diffusion model. Tan et al. (Tan et al., 2023) adopt a pre-trained CNN model as transformation models convert images into gradients. Ojha et al. (Ojha et al., 2023) proposed based on pre-trained vision-to-language models to develop a universal generated image detector. In orthogonal, a large amount of work focuses on the detection of Deepfake images (Guarnera et al., 2020; Li et al., 2020; Zhao et al., 2021a; Shiohara & Yamasaki, 2022), such as artificially creating general forged boundaries (Shiohara & Yamasaki, 2022), detect shifts in the frequency domain (Guo et al., 2023), and detect Deepfake from inconsistencies in face image regions (Zhao et al., 2021b).
So far, there are some studies on detecting Deepfake videos(Güera & Delp, 2018; Masi et al., 2020; Agarwal et al., 2020; Haliassos et al., 2021; Wang et al., 2023c), but there is a lack of research specifically focused on detecting generated videos. We hope this paper can bring exploratory and insightful contributions to this domain of study.
3 Problem Formulation
Generated Video Detection. In addition to the generalization ability of generated video detectors to unseen generation models, we also focus on the generalization to cross-category videos. So, the generated video detection task is defined as follows:
T2V-zero Show1 MScope ZScope Real
Time 0 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Time 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: We generated four sub-datasets based on the prompts of real videos and generation models based on four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and SHOW-1. Each pair of positive and negative samples follows the same prompt, and we show the first five frames of four pairs of positive and negative samples corresponding to one of the prompts: A person is cutting a carrot.
Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refers to each real video and its corresponding content prompt. And let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denotes video generation models, where $g_{j}$, $j\in[1,m]$ represents a specific video generation model. Accordingly, based on the prompt $p_{i}$ of each real video $r_{i}$, it generates $m$ corresponding generated videos, resulting in a set of fake video collections $F_{j}\in F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$. the generated video detection dataset $D_{j}$in $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $x\in\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$and $y\in\mathcal{Y}=\{0,1\}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
The goal of generated video detection is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generate model $m_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and achieve generalization on unseen generate models $g{k}$, where $k\in[1,m],k\neq j$. This can be achieved by minimizing classification error of $\mathcal{F}$ on training data $D^{*}_{i}$:
$\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{(x,y)\in D^{*}_{i}}\mathcal{L}(f(x),y)$ (1)
Where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$ We implement the evaluation of detection of cross-category generated videos by ensuring that the training set and the test set do not have videos with the same prompt.
4 Dataset Construction
To benchmark generated video detectors, we created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ dataset. In this section, we mainly introduce the construction of the data set and conduct a comprehensive evaluation of the dataset.
4.1 Dataset Composition
The $GeneratedVideoForensics$ (GVF) dataset $\mathcal{D}$ consists of 964 triples and can be represented as $\mathcal{D}=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $F$ represents multiple fake videos generated by different generation models based on the text prompt $P$. In order to examine the generalization of the detectors, we selected four open source generate models with different generation methods and model frameworks: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a), ZeroScope (Sterling., June 2023) and SHOW-1 (Zhang et al., 2023) as generation models for negative samples. Therefore, the GVF dataset could be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show1. Figure  3 shows an example of one of the triples.
Figure 4: Data distribution over categories under the “major content (upper) and “attribute control” (lower) aspects.
Prompt. To approximate the distribution of videos in real task scenarios, we collected 964 prompts and corresponding real videos from existing open-domain text-to-video datasets: MSVD (Chen & Dolan, 2011) and MSR-VTT (Xu et al., 2016) datasets. Following the FTEV (Liu et al., 2023), we consider the collection of prompts from the following four aspects to simulate the sample distribution in real scenarios: the spatial and temporal major content it describes and the spatial and temporal attributes it controls during video generation. Spatial content is classified as people, animals, plants, food, vehicles, buildings, and artifacts, with corresponding spatial attributes of quantity, color, and camera view. Temporal content and attributes encompass actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order. This ensures that our data set is comprehensive and balanced. Figure  4 shows the data distribution on categories under different aspects, which shows that the GVF dataset contains enough data to cover different category ranges.
Figure 5: Manual evaluation results across spatial major contents.
4.2 Dataset Evaluation
Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation ideas and model architectures, to restore better the challenges posed by invisible generation models. Figure 5 shows the performance of videos generated by the above four generation models on different attributes under the FETV  (Liu et al., 2023) standard. In order to generate different kinds of content, it is obvious to note that each of these models has its advantages and disadvantages. Text2Video-zero, for example, has the best impact on the spatial quality of the video but the worst impact on the temporal continuity. On the contrary, ZeroScope has a better effect on temporal continuity but has the worst effect on spatial quality. The choice of these video generation models, each possessing distinct advantages and disadvantages, ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
5 Methodology
5.1 Probe experiments of STCNN models
Initially, we considered using a spatiotemporal convolutional model (STCNN) to train the generally generated video detector. However, we discovered that these models are only effective on the specific generative models they were trained on and needed more generalizability, failing to discern unseen T2V generation models.
To analyze what information the STCNN network was learning during discrimination, or rather, which features within the video it was utilizing to distinguish between authentic and fabricated content, we designed a sophisticated probe experiment with two novel datasets under the same model framework: The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to see if the model can still differentiate based solely on the temporal discontinuities of the videos. The second dataset was created by randomly selecting a frame from the current test set and replicating it multiple times to see if the model could still differentiate based on the spatial artifacts of the video frame alone.
The experimental results in table 1 led us to the following conclusion: We found that the model’s attention to spatial artifacts and temporal artifacts was distinct, and for developing a generalized discriminator, they did not play the same role. Therefore, we attempt to build detectors based on spatial artifacts and temporal artifacts separately, individually verifying the role each played in the learning process of the universally generated video detector.
Table 1: Spatiotemporal Convolutional Neural Networks (3DCNN) rely on easily detectable spatial artifacts.
Method Artifacts Difference
Temporal Spatial
ACC AP ACC AP ACC AP
I3D 51.17 51.61 86.67 98.61 35.50 47.00
Slow 50.04 50.37 93.81 99.51 43.77 49.14
X3D 50.04 54.31 78.39 92.84 28.35 38.53
Mvit 50.48 55.29 72.16 92.59 21.68 37.30
5.2 Shortcomings of learning from spatial artifacts
To analyze the role of spatial artifacts, we first work with existing T2I models which have been used as image-level detectors for generated videos, and we test them on the GVF dataset. As shown in Table 2, these detectors demonstrate superior capabilities in capturing the spatial artifacts of specific generation models. However, they lack generalization to unseen generation models.
To find out why these detectors lack generalization to unseen generate models, we visualized the feature space of the detector (Wang et al., 2020) and examined the distribution of four sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, consisting of fake videos from the sub-dataset Text2Video-zero, (iii) Show1, composed of fake videos from the sub-dataset Show-1, and (iv) MP, comprised of fake videos from the sub-dataset ModelScope, (v) ZP, comprised of fake videos from the sub-dataset ZeroScope. For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE (Van der Maaten & Hinton, 2008).
As shown in Figure 6, on the one hand, the detector has ideally classified the real video and the fake video generated by the specific generate model, which means that the detector has fully captured the spatial artifacts of the specific generate model Text2Video-zero  (Khachatryan et al., 2023). On the other hand, the detector cannot distinguish fake videos generated by unseen models from real videos. Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting other video generation models? We illustrate this by visualizing the frequency spectrum of video frames generated by various video generate models. As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent; capturing more general spatial artifacts appears to be extremely difficult.
Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models. The feature space used is one trained to distinguish fake videos (Text2Video-zero) from real videos.
Refer to caption Real
Refer to caption T2V
Refer to caption Show1
Refer to caption MP
Refer to caption ZP
Figure 7: Average spectrum of video frames for real videos and fake videos generated by different video generation models.
Figure 8: Overview of the DeCoF framework. We first get real video and generated video features by the pre-trained model CLIP:VIT, which is designed to eliminate the impact of spatial artifacts on capture temporal artifacts, and then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of real and fake video.
5.3 Eliminate impact of spatial artifacts
There is no doubt that real video frames are continuous, and we can even predict any frame of the video through other frames. However, mismatches are prone to occur between the generated video frames, resulting in the inability to meet the continuity of the video frames, that is, temporal artifacts of the generated video. Given a set of video frames $V=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a particular function that measures the continuity between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video.
After analyzing the results of the above exploration, it appears that the detection model tends to learn spatial artifacts that are easy to capture but do not have generalization properties. The difficulty lies in how to eliminate the impact of spatial artifacts on capturing temporal artifacts.
Intuitively, if we can find a particular mapping function $g$ with the following properties: The distance between frame mappings is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts. In fact, we choose ViT-L/14  (Dosovitskiy et al., 2020) trained for the task of image-language alignment as mapping function $g$, then by mapping the video frames $F_{i}$ to $g(F_{i})$, the decoupling of temporal artifacts and spatial artifacts can be achieved. As a result, the original problem can be expressed as $f(g(F_{1}),g(F_{2}),\cdots,g(F_{L}))=y$.
Model framework. In order to learn video frame consistency for capturing temporal artifacts, we design our detection with the Consistency of video Frames (DeCoF), in which the critical verification module consists of two transformer layers and an MLP head, As shown in Figure  8. More specifically, given a video input $V$, The frozen parameter image encoder (ViT-L/14) will extract the features of the video frames and generate a feature sequence $S=\{g(F_{1}),g(F_{2}),\cdots,g(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature. Next, the position encoding and a learnable class embedding representing the classification output are prepended to the sequence $S$ and then passed through a sequential module composed of two layers of transformers. Finally, a classification header is appended to the transformer layer, resulting in a probability distribution over the target class, and we train our model via a simple cross-entropy loss.
In addition to decoupling temporal and spatial artifacts, our approach significantly reduces computational complexity and memory requirements since we only need to learn anomalies between video frame features. Overall, our approach is simple yet effective.
6 Experiments
In this section, we introduce the experiment setup and provide extensive experimental results to demonstrate the superiority of our approach.
Table 2: Comprehensive comparison of our method with previous detectors and spatiotemporal convolutional networks, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model. * Denotes our retraining on a subset of generated video forensics with official code. We report ACC (%) and AP (%) (ACC and AP in the Table).
Method Training dataset Generate mode Detection Level Tested generate video models Total Avg.
Text2Video-zero Show1 ModelScope ZeroScope
ACC AP ACC AP ACC AP ACC AP ACC AP
CNNDet(0.5) LSUN ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
CNNDet(0.1) LSUN ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
DIF SDV2 Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
DIRE LSUN-B Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
Lgard LSUN ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
CNNDet (0.5)* Text2Video-zero T2V Image 98.64 99.97 51.93 61.64 51.54 68.91 54.96 72.51 64.27 75.76
CNNDet (0.1)* Text2Video-zero T2V Image 96.84 99.99 53.93 60.99 55.34 68.87 60.69 73.80 66.70 75.91
Lgard* Text2Video-zero T2V Image 98.45 100.00 56.19 78.91 51.03 66.30 54.64 75.93 65.08 80.29
DIRE* Text2Video-zero T2V Image 99.17 99.99 49.93 56.74 52.83 77.58 51.28 67.12 63.30 75.36
I3D* Text2Video-zero T2V Video 98.00 99.87 49.75 70.41 49.25 71.43 50.50 51.85 61.88 73.39
Slow* Text2Video-zero T2V Video 93.30 99.48 50.48 52.09 50.59 49.43 49.41 37.29 60.95 59.57
X3D* Text2Video-zero T2V Video 94.84 98.61 49.41 33.52 48.35 43.07 49.85 30.89 60.61 51.52
Mvit* Text2Video-zero T2V Video 98.42 99.97 49.96 45.93 49.49 38.98 51.68 32.60 62.39 54.37
DeCoF Text2Video-zero T2V Video 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
DeCoF Show1 T2V Video 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
DeCoF ModelScope T2V Video 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
DeCoF ZeroScope T2V Video 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
6.1 Experimental Setup
Data pre-processing.We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalization of the detector. We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
Evaluation metrics. Following previous methods  (Wang et al., 2020, 2023b; Frank et al., 2020), we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors. The threshold for computing accuracy is set to 0.5 following  (Wang et al., 2020, 2023b). For image-based methods, we average frame-level predictions into corresponding video-level predictions.
Baseline.  (i) CNNDet (Wang et al., 2020) proposes a model for detecting images generated by CNN. Through a large amount of data enhancement, the model has generalization on various images generated by CNN.  (ii) DIF  (Sinitsa & Fried, 2023) proposes a few-shot learning method for image synthesis.  (iii) DIRE (Wang et al., 2023b) proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models.  (iv) Lgard (Tan et al., 2023) used a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts.  (v) I3D (Carreira & Zisserman, 2017) introduce a new Inflated 3D ConvNet that is based on 2D ConvNet inflation.  (vi) Slow (Feichtenhofer et al., 2019) proposed a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames.  (vii) X3D (Feichtenhofer, 2020) designed a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy.  (viii) Mvit  (Fan et al., 2021) combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition.
6.2 Detailed Evaluation
We conduct comprehensive and challenging experiments to answer the following questions:
Figure 9: Robustness to unseen perturbations. The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression. We test the detectors trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
Q1. How does our detector perform at detecting unseen T2V models compared to previous detectors?
As mentioned in Section 5.2, we found that previous detectors exhibit extraordinary capabilities when detecting specific generative models but lack generalization when unseen generative models are present. In the table 2, we quantitatively evaluate CNNDet (Wang et al., 2020), DIF (Sinitsa & Fried, 2023), DIRE (Wang et al., 2023b), Lgard (Tan et al., 2023), using the pre-trained weights provided by the official repository. Although these detectors have surprising performance in detecting generated images, their performance drops significantly on the task of detecting generated videos, with both ACC and AP below 60%. As a result of retraining the above four detectors on the sub-dataset Text2Video-zero, their performance on specific video generation models greatly improved, but they still performed poorly for unseen generation models, and ACC and AP scores are still lower than 60%, and even the retrained spatiotemporal convolutional network displays similar results. On the contrary, compared with previous detectors, our generated video detection method has excellent generalization for various unseen generation models. Specifically, on the task of generating video detection, the ACC increased by 17.74% $\sim$ 31.38% and the AP by 16.73% $\sim$ 49.18%.
We also conducted cross-dataset evaluations. Table  2 shows that on the task of generating video detection, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%. This is because Text2Video-zero has the worst temporal continuity in all four T2V generation models.
Table 3: Performance of DeCoF on commercial generation models.
Training dataset Tested generate video models Total Avg.
Gen-2 Pika
ACC AP ACC AP ACC AP
Text2Video-zero 99.49 100.00 99.49 99.98 99.49 99.99
Show1 98.46 99.96 98.46 99.85 98.46 99.91
ModelScope 98.97 99.99 96.89 99.73 97.93 99.86
ZeroScope 98.97 99.99 94.84 99.52 96.91 99.76
Q2. How does our detector maintain robustness against noise and disturbances?
In practical detection scenarios, the robustness of the detector to unseen perturbations is also crucial. Here, we mainly focus on the impact of two perturbations on the detector: Gaussian blur and JPEG compression. The perturbations are added under three levels for Gaussian blur ($\sigma$ = 1, 2, 3) and five levels for JPEG compression (quality = 90, 80, 70, 60, 50). Figure  9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations. Although the model’s performance in the face of perturbations is inconsistent due to differences in training data sets, Our method remains effective under any degree of arbitrary perturbation, and the AP of models trained on any sub-dataset is greater than 80%.
Q3. How does our detector perform on commercial generation models?
To further evaluate the effectiveness of our detection method, we constructed two additional test datasets using two well-known commercial proprietary T2V generation models, Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023). Based on the datasets generated by the four open-source T2V models shown in Table  3, the detector was trained on Text2Video-zero, Show1, ModelScope, and ZeroScope. When confronted with the commercial generation model, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2, demonstrating the strong generalization capability of our model DeCoF for unseen T2V models. Even when trained on specific T2V models, the model is highly accurate in detecting newly generated video.
7 Conclusions and Future Work
The paper presents a groundbreaking investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content. We construct the first public dataset for real and fake video detection, providing a solid benchmark for community research. At the same time, we propose a simple yet effective detection model, which has proven its strong generalizability and robustness on multiple open-source and commercial closed-source models. We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
8 Impact Statements
Our research focuses on using Machine Learning to detect generated videos. We developed the DeCoF model and shared our dataset to support detecting synthetic videos. These works are crucial to protect digital content and prevent the spread of disinformation. Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology. We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity. We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
Appendix A $GeneratedVideoForensics$ Dataset
A.1 Dataset Composition
The following section provides a more detailed description of how the generated video forensics dataset was constructed and evaluated. The entire GVF dataset is generated in two steps:
1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  (Chen & Dolan, 2011) and MSR-VTT  (Xu et al., 2016) datasets. When selecting samples, we followed the work of FETV  (Liu et al., 2023), Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios. More specifically, in the work of FETV, they classify each prompt controlling video generation from two aspects: the main content it describes and the attributes it controls during video generation. The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations. The corresponding spatial attribute controls are divided into quantity, color, and camera view. Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
2) Selection of generation models: Although there are many high-quality generative models on the market, we do not want to choose the best. On the contrary, we have selected four generative models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of Generative models appear, how do we deal with the problems caused by invisible generative models? In the end, the four most representative generative models were selected as our generative models: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a),ZeroScope(Sterling., June 2023),Show1 (Zhang et al., 2023). At the same time, we also selected the two most popular models as the secret test dataset: Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023).
If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$. If divided by the generation model, the GenerateVideoForensics data set can be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and SHOW-1. Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
A.2 Dataset Evaluation
As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV (Liu et al., 2023).
Table 4: Some examples of prompts. Each prompt is composed of a random combination of four attributes: main content of time, main attributes of time, main content of space, and main attributes of space. Here we list some examples of prompts, red for content and blue for attributes.
Main content Attribute
control
Example
Spatial People Quantity four friends are driving in the car
Animals Color a white bird chirps and plays on a handmade staircase
Plants Camera View a beautiful view of greenery and river from a top view
Food Color in cup full of chocolate then we add yellow cream to it so yummy
Temporal Actions Event Order a red-headed woman is shown walking and then sitting
Actions Speed a man fastly cross the road by run and get in to the car
Kinetic Motions Motion Direction the big black tractor is moving away from it s carrier
Fluid Motions Speed a bucket of slime is quickly poured on top of a woman’s head
Figure 10: Data distribution over categories under the “major content” (upper) and “attribute control” (lower) aspects.
Attribute statistics of prompts. From the four aspects mentioned before, the main content of space and time and the main attributes of space and time, we provide attribute statistics of prompts of the data set. As the figure 10 shows, the various attributes of prompts are different because one prompt may have multiple attributes at the same time. At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different. You may ask why time main content and time attributes are so rare. In fact, not every prompt in the MSVD and MSR-VTT datasets has a time description, and we only ensure the existence of this type of prompt in the data set. Here, we also provide some specific examples of prompts, As shown in the table 4.
Evaluation of different generation models. Here, we present our assessment of four generation models according to the FETV standard (Liu et al., 2023). We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames. (ii) Temporal quality focuses on the temporal coherence of video frames. (iii) Overall alignment measures the overall alignment between a video and the given text prompt. (iv) Fine-grained alignment measures the video-text alignment regarding specific attributes.
As shown in figure 11, Regarding spatial categories, the performance quality of these four data sets varies, and the performance of the same data set differs significantly between the static quality and temporal quality levels. The videos that include people, animals, and vehicles have poor spatial quality, but the videos that include plants, scenery, and natural items have greater spatial quality. The reason behind this is that the videos in the first three categories contain numerous high-frequency intricacies, such as face and finger, which are more difficult to comprehend compared to the low-frequency information found in the final two categories, such as water and clouds.
In terms of static quality, Text2Video-zero performs well but performs poorly in terms of temporal quality. Text2Video-zero inherited the ability to generate images from Stable Diffusion, but it lacks temporal coherence in its videos due to the lack of video training. SHOW-1 is second only to Text2Video-zero in terms of static quality, and its temporal quality is also superior. In comparison to ModelScopeT2V, ZeroScope is inferior to ModelScope in temporal quality but performs worse in static quality. The reason for this may be that ZeroScope has been primarily designed for videos with a 16:9 aspect ratio and without watermarks, rather than focusing on improving visual quality and aligning video text.
It is important to note that while the four models have better control over color and camera view, they do not perform as well when it comes to accurately controlling attributes such as quantity, motion direction, and order of the events. In terms of speed and motion direction attribute control, ModelScope and ZeroScope are particularly promising, whereas Text2Video-zero demonstrates poor performance.
On different attributes of the four models, fine alignment and global alignment perform similarly, although there are some exceptions. This is due to the fact that other properties may have an effect on the overall alignment. Therefore, it is more accurate to assess controllability using fine alignment when examining a specific attribute, whereas global alignment is a reasonable approximation.
Appendix B Parameter Description
In this section, we provide a detailed description of various parameters of the experiment, including the hyperparameters of the model and the details of its training.
Training details. During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets. The numbers of the training set, validation set, and test set are 771, 97, and 96, respectively. We freeze the parameters of CLIP:ViT (Dosovitskiy et al., 2020), making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9.
Data pre-processing. All the following experiments are performed on our GVF dataset. In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 (Zhang et al., 2023), ModelScope (Wang et al., 2023a) and ZeroScope (Sterling., June 2023) Since the resolution and length of the videos generated by the four generative models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224. During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224. Some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
Model hyperparameters. Our model consists of two layers of transformer and an MLP head is an adaptation of the sequence to sequence trans- former, where the dimension and feature dimension of MLP is 768, path_size is 32, the dropout parameter is 0.1, and the number of attention heads is 4.
Table 5: A comparison of the impact of different mapping functions.
Mapping Function Pretrain Tasks Tested generate video models Total
Avg.
Text2Video-zero Show1 ModelScope ZeroScopre
ACC AP ACC AP ACC AP ACC AP ACC AP
ViT-L/14 CLIP 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
Resnet50 CLIP 96.92 99.82 52.60 80.66 57.22 80.19 50.00 67.65 64.19 82.08
ViT-B/16 ImageNet
Resnet50 ImageNet
Figure 11: The manual evaluation provided results across major static and temporal contents and attributes to control video-text alignment.
Appendix C Different Mapping Functions
In this section, we examine the impact of different mapping functions.
The mapping function should have two properties. One is that the distance between features in the feature space is inversely proportional to the similarity between images, and the other is that it is insensitive to spatial artifacts. The second requirement is easy to implement as long as the mapping function is not used to train the fake content detection task. For the first requirement, we explored several different mapping functions, as shown in the table 5.
First, we discussed the two different tasks of ImageNet and CLIP. Interestingly, the encoder trained on ImageNet caused this problem not to converge. We considered two reasons: one is the size of the training data, and the other is Due to the difference in the task itself; we do not want the model to classify images into specific items or biological categories but rather cluster similar images. At the same time, we compared the impact of different model frameworks. Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function. Overall, the mapping function largely determines the effectiveness of our approach.
Appendix D More Visual Examples
Here, we provide more examples of generated videos. Figures 12 to Figures 15 are examples of real videos and generated videos corresponding to prompts.
Figure 12: Real videos and videos generated by four generation models according to prompt: A person mixing ingredients in a mixing bowl.
Figure 13: Real videos and videos generated by four generation models according to prompt: A kitten moves about.
Figure 14: Real videos and videos generated by four generation models according to prompt: A young man is seated and playing a guitar.
Figure 15: Real videos and videos generated by four generation models according to prompt: A boy rides a bike on a fence.

### v3

@@
-DeCoF: Generated Video Detection via Frame Consistency
+DeCoF: Generated Video Detection via Frame Consistency: The First Benchmark Dataset
 Abstract
-The escalating quality of video generated by advanced video generation methods leads to new security challenges in society, which makes generated video detection an urgent research priority.
-To foster collaborative research in this area, we construct the first open-source dataset explicitly for generated video detection, providing a valuable resource for the community to benchmark and improve detection methodologies.
-Through a series of carefully designed probe experiments, our study explores the significance of temporal and spatial artifacts in developing general and robust detectors for generated video.
-Based on the principle of video frame consistency, we introduce a simple yet effective detection model (DeCoF) that eliminates the impact of spatial artifacts during generalizing feature learning.
-Our extensive experiments demonstrate the efficacy of DeCoF in detecting videos produced by unseen video generation models and confirm its powerful generalization capabilities across several commercial proprietary models.
+The escalating quality of video generated by advanced video generation methods results in new security challenges, while there have been few relevant research efforts: 1) There is no open-source dataset for generated video detection, 2) No generated video detection method has been proposed so far.
+To this end, we propose an open-source dataset and a detection method for generated video for the first time.
+First, we propose a scalable dataset consisting of 964 prompts, covering various forgery targets, scenes, behaviors, and actions, as well as various generation models with different architectures and generation methods, including the most popular commercial models like OpenAI’s Sora and Google’s Veo.
+Second, we found via probing experiments that spatial artifact-based detectors lack generalizability.
+Hence, we propose a simple yet effective detection model based on frame consistency (DeCoF), which focuses on temporal artifacts by eliminating the impact of spatial artifacts during feature learning.
+Extensive experiments demonstrate the efficacy of DeCoF in detecting videos generated by unseen video generation models and confirm its powerful generalizability across several commercially proprietary models.
+Our code and dataset will be released at https://anonymous.4open.science/r/DeCoF-8394.
 1 Introduction
-Following the success of diffusion models in generating highly realistic images (Nichol & Dhariwal, 2021; Rombach et al., 2022; Parmar et al., 2023), a growing number of researchers have shifted their focus to the advancement of these models for video generation (Kawar et al., 2023; Blattmann et al., 2023; Zhang et al., 2023).
-Such developments have turned diffusion models into robust and versatile video generators, enabling the creation of high-quality videos from general conditional inputs, such as text and images.
-While this powerful and flexible generation capacity has the potential to alleviate the stress of daily life and inspire creativity in work, it concurrently introduces a new crisis of trust (Sharma et al., 2023; Barrett et al., 2023).
-Therefore, there is an urgent need for detectors capable of identifying generated videos.
-In this paper, we propose a challenging task: Generated Video Detection, which aims to create universal detectors to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models.
-To evaluate generated video detectors, we create the first comprehensive generated video detection dataset: the $GeneratedVideoForensics$ (GVF) dataset.
-The GVF dataset encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments.
-Moreover, pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos.
-Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures (Khachatryan et al., 2023; Zhang et al., 2023; Wang et al., 2023a; Sterling., June 2023), such as pixel-space or latent-space models, which aims to measure the generality of the generated video detectors.
-Figure 1: Can the detector trained on a specific video generation model cope with the challenges presented by unseen generation models?
-Considering that video content contains individual frame contents and smooth consistency across multiple frames, a natural idea is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) models to achieve generated video detection.
-Our exploratory experiments found that these models have excellent discriminative capabilities for specific T2V models but cannot be applied to detect unseen T2V models.
-Refer to caption (a) Geometry
-Refer to caption (b) Layout
-Refer to caption (c) Frequency
-Refer to caption (d) Color
+Following the success of diffusion models in generating highly realistic images [28, 34, 30], increasing attention has been paid to video generation [21, 4, 51], enabling the creation of high-quality videos from general conditional inputs such as text and images.
+However, this powerful generation capacity also introduces a new crisis of trust [36, 3].
+Therefore, there is an urgent need for detectors for generated videos.
+Figure 1: Can the detector trained on a specific video generation model detect videos generated by unseen generation models?
+In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models.
+Moreover, we propose the first comprehensive generated video detection dataset, the $GeneratedVideoForensics$ (GVF) dataset, for evaluating the performances of generated video detectors.
+GVF encompasses a wide range of scene contents and motion variations, approximating the authenticity verification problem caused by the T2V generation model in diverse actual environments.
+Pairs of real and generated videos in the dataset are constructed using the same prompt to ensure that the generated video detectors learn subtle distinctions to identify the authenticity of videos rather than learning differences between the content of the real and fake videos.
+Furthermore, we construct generated samples corresponding to real videos based on various T2V models with different generation methods and model architectures, such as pixel-space or latent-space models [22, 51, 46, 42], to measure the generalizability of the generated video detectors.
+(a) Geometry
+(b) Layout
+(c) Frequency
+(d) Color
 (e) Temporal
-Figure 2:  Spatial and temporal artifacts on generated video.
-We illustrate the spatial and temporal artifacts present in the generated videos.
+Figure 2: Illustration of spatial and temporal artifacts on generated video.
 Spatial artifacts: (a) errors in geometric appearance, (b) errors in image layout, (c) frequency inconsistency, (d) color mismatch; Temporal artifacts: (e) mismatch between frames.
-Further probe experiments in Table 1 reveal that the spatiotemporal convolutional models predominantly identify video authenticity by detecting spatial artifacts, resulting in generated video detection that has degenerated into a 2D problem.
-A critical question arises: Spatial and temporal artifacts, which artifacts are key to constructing universal detectors?
-To explore this problem, we first analyze the possibility of building a universal generated video detector through spatial artifacts.
-We convert the 3D problem of generated video detection into a 2D problem, generated image detection, and implement T2V detection based on existing text-to-image (T2I) detection methods.
-In section  5.2, we evaluate the performance of these generated image detectors on generated video detection, noting that the reason why these detectors lack generality to unseen generative models is that capturing specific spatial artifacts to identify video authenticity, while spatial artifacts are specific to the T2V model, these detectors are unable to capture more general and high-level spatial artifacts to establish better decision boundaries.
-So far, we conclude that spatiotemporal convolutional models primarily identify video authenticity by detecting spatial artifacts but are unable to detect more general spatial artifacts.
-Since spatial artifacts are easily detected, learning of temporal artifacts is difficult to perform.
-To rectify this biased learning approach, mitigate the adverse effects induced by spatial artifacts, and construct a more generalizable generated video detector by capturing temporal artifacts, we propose to decouple temporal and spatial artifacts by mapping video frames to feature space, where the distance between features is negatively correlated with image similarity, through a pre-trained CLIP:ViT (Dosovitskiy et al., 2020).
-More importantly, the features of real video frames are still continuous after being mapped by the mapping function with the above requirements.
-Meanwhile, the frames of the generated video with inter-frame mismatch will have anomalies in their features.
-Besides, it greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned.
-Specifically, we only implement the learning of inter-feature anomalies through a temporal artifact modeling module with two transformer layers (Vaswani et al., 2017).
-Extensive exploring experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection, especially significant generalization capabilities for unseen generation models.
-Our contributions are summarized as follows:
-  • We present the first dataset that is comprehensive and scalable as new models emerge for benchmarking generated video detectors.
-  • We explore the possibility of detecting generated videos by capturing spatial and temporal artifacts and point out the drawbacks of the detector by capturing spatial artifacts and their causes.
-  • We develop a universal generated video detector named DeCof, validating through extensive experiments, including tests on commercial proprietary models.
+Considering that authentic videos should present smooth inter-frame consistency, a naive approach is to utilize end-to-end spatiotemporal convolutional neural networks (STCNN) for generated video detection.
+However, we found via exploratory experiments that these models lack generalizability to detect videos generated by unseen T2V models.
+Further probe experiments revealed that spatiotemporal convolutional models primarily recognize video authenticity through spatial artifacts (Figure 2 illustrates two types of artifacts in generated videos), degrading video detection into a 2D problem.
+By analyzing the performance of existing text-to-image (T2I) detection works [48, 43, 47], we found it difficult to capture general spatial artifacts, as the contrast and style of video frames generated by different generation models vary.
+Therefore, to develop a universal generated video detector, we propose the first temporal-artifact-based detection method.
+More specifically, we decouple temporal and spatial artifacts by mapping video frames to feature space where the inter-feature distance is negatively correlated with image similarity, to detect the anomalies caused by inter-frame inconsistency.
+Our method greatly reduces computational complexity and memory requirements since only anomalies between features need to be learned.
+Extensive experiments demonstrate that our method exhibits extraordinary capabilities in generated video detection and excellent generalizability for unseen generation models, even on the latest generation models Sora [41] and Veo [45].
+Our contributions can be summarized as follows:
+  • We present the first comprehensive and scalable dataset for benchmarking generated video detectors.
+  • Through a series of probing experiments, we have demonstrated the flaws in detecting generated videos from spatial artifacts.
+  • We develop the first temporal-artifact-based universal generated video detector named DeCof and validate its performance through extensive experiments, including tests on the latest commercial proprietary models.
 2 Related work
 2.1 Generation Models
-Recently, with the success of diffusion models in generating images (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021; Esser et al., 2021; Rombach et al., 2022; Peebles & Xie, 2023), a number of Research explores text-to-video Diffusion Models (VDMs).
-Some work (Ramesh et al., 2022; Blattmann et al., 2023; Khachatryan et al., 2023; Wang et al., 2023a)ground their models on latent-based VDMs, for example, Text2Video-zero (Khachatryan et al., 2023) generates videos by using motion dynamics to enhance the latent code of a text-to-image stable diffusion model without requiring training on any video data, and ModelScopeT2V (Wang et al., 2023a) extends it with T2I stable diffusion model of convolutional and attentional blocks and trained using image-text and video-text datasets, ZeroScope (Sterling., June 2023) is a watermark-free T2V generative model trained with the weights of ModelScopeT2V, which is optimized for 16:9 compositions.
-On the other hand, PYoCo (Ge et al., 2023)and Make-A-Video (Singer et al., 2022) anchor their models on pixel-based VDMs.
-Show1 (Zhang et al., 2023) integrates pixel-based and latent-based VDMs, using Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, while latent-based VDMs upscale these low-resolution videos from pixel-based VDMs to then create high-resolution videos with low computation cost.
-2.2 Generated Content Detections
-In recent years, a large amount of work has focused on generated image detection (McCloskey & Albright, 2019; Zhang et al., 2019; Barni et al., 2020; Corvi et al., 2023; Sha et al., 2023), especially for unseen generation models.
+Recently, with the success of diffusion models in generating images [20, 40, 28, 9, 34, 31], a number of research explores text-to-video Diffusion Models (VDMs).
+Some works [33, 4, 22, 46] ground their models on latent-based VDMs, e.g., Text2Video-zero [22] generates videos by using motion dynamics to enhance the latent code of text-to-image(T2I) model [34] without the need for training on any video data, and ModelScopeT2V [46] extends T2I model with temporal convolution and attention blocks trained with image-text and video-text datasets.
+ZeroScope [42] is a watermark free version of ModelScopeT2V, and has been optimized for 16:9 composition.
+On the other hand, PYoCo [14] and Make-A-Video [38] anchor their models on pixel-based VDMs.
+Show1 [51] integrates Pixel-based VDMs to produce videos of lower resolution with better text-video alignment, then upscale them with latent-based VDMs to create high-resolution videos with low computation cost.
+T2V-zero
+Show1
+MScope
+ZScope
+Real
+Time 0
+Time 1
+Time 2
+Time 3
+Time 4
+Figure 3: We generated four sub-datasets based on the real videos using the prompts and selected generation models employing four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1.
+Each pair of positive and negative samples use the same prompt.
+Here shows the first five frames of four sample pairs using the prompt: A person is cutting a carrot.
+2.2 Generated Content Detection
+Many studies have focused on generated image detection [27, 52, 2, 7, 35], especially for unseen generation models.
 A simple classifier trained by Wang et al.
-(Wang et al., 2020) on pagan-generated images using multiple data augmentations can be generalized to 20 unseen other GAN-generated images.
+[47] on ProGAN-generated images can be generalized to 20 unseen other GAN-generated images.
 Frank et al.
-(Frank et al., 2020) concluded that the artifacts of the generated images are caused by the upsampling method.
+[13] conclude that the artifacts of the generated images are caused by the upsampling method.
 Wang et al.
-(Wang et al., 2023b) identify generated images through the difference of images reconstructed by a pre-trained diffusion model.
-Tan et al.
-(Tan et al., 2023) adopt a pre-trained CNN model as transformation models convert images into gradients.
+[48] identify generated images through the difference of images reconstructed by a pre-trained diffusion model.
 Ojha et al.
-(Ojha et al., 2023) proposed based on pre-trained vision-to-language models to develop a universal generated image detector.
-In orthogonal, a large amount of work focuses on the detection of Deepfake images (Guarnera et al., 2020; Li et al., 2020; Zhao et al., 2021a; Shiohara & Yamasaki, 2022), such as artificially creating general forged boundaries (Shiohara & Yamasaki, 2022), detect shifts in the frequency domain (Guo et al., 2023), and detect Deepfake from inconsistencies in face image regions (Zhao et al., 2021b).
-So far, there are some studies on detecting Deepfake videos(Güera & Delp, 2018; Masi et al., 2020; Agarwal et al., 2020; Haliassos et al., 2021; Wang et al., 2023c), but there is a lack of research specifically focused on detecting generated videos.
-We hope this paper can bring exploratory and insightful contributions to this domain of study.
+[29] propose to develop a universal generated image detector based on pre-trained vision-to-language models.
+Meanwhile, a large amount of work focuses on Deepfake image detection [16, 23, 53, 37], using methods such as artificially creating general forged boundaries [37], detecting shifts in the frequency domain [18], and detecting inconsistencies in face image regions [54].
+So far, there are some studies on detecting Deepfake videos [17, 26, 1, 19, 49], but research on detecting generated videos remain unexplored.
+This paper targets at bringing exploratory and leverageable contributions to this new direction.
 3 Problem Formulation
 Generated Video Detection.
-In addition to the generalization ability of generated video detectors to unseen generation models, we also focus on the generalization to cross-category videos.
-So, the generated video detection task is defined as follows:
-T2V-zero Show1 MScope ZScope Real
-Time 0 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Time 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Time 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Time 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Time 4 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 3: We generated four sub-datasets based on the prompts of real videos and generation models based on four different architectures and generation methods: Text2Video-zero, ModelScopeT2V, ZeroScope, and SHOW-1.
-Each pair of positive and negative samples follows the same prompt, and we show the first five frames of four pairs of positive and negative samples corresponding to one of the prompts: A person is cutting a carrot.
-Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refers to each real video and its corresponding content prompt.
-And let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denotes video generation models, where $g_{j}$, $j\in[1,m]$ represents a specific video generation model.
-Accordingly, based on the prompt $p_{i}$ of each real video $r_{i}$, it generates $m$ corresponding generated videos, resulting in a set of fake video collections $F_{j}\in F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$.
-the generated video detection dataset $D_{j}$in $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $x\in\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$and $y\in\mathcal{Y}=\{0,1\}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
-The goal of generated video detection is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generate model $m_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and achieve generalization on unseen generate models $g{k}$, where $k\in[1,m],k\neq j$.
+In addition to the generalizability to unseen generation models, we also take the generalizability to cross-category videos into account.
+The generated video detection task can be defined as follows:
+Given a collection of real videos denoted as $R=\{r_{1},r_{2},\cdots,r_{n}\}$, a set of content prompts $P=\{p_{1},p_{2},\cdots,p_{n}\}$, where $r_{i}$ and $p_{i}$, $i\in[1,n]$ refer to each real video and its corresponding content prompt.
+Let $G=\{g_{1},g_{2},\cdots,g_{m}\}$ denote a set of video generation models.
+Using the prompt $p_{i}$ of each real video $r_{i}$, $G$ generates $m$ corresponding videos, resulting in a set of fake video collections $F=\{F_{1},F_{2},\cdots,F_{m}\}$, where $F_{j}=\{g_{j}(p_{1}),g_{j}(p_{2}),\cdots,g_{j}(p_{n})\}$, $j\in[1,m]$.
+As such, we form the generated video detection dataset $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{m}\}$, where $D_{j}=\{(x,y)^{(i)}\}_{i=1}^{n}$, with $\mathcal{X}=\{R,F_{j}\}\subset\mathbb{R}^{L*H*W*C}$, $\mathcal{Y}=\{0,1\}$, $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ denoting a video and its label, $L,H,W$ and $C$ correspond to the video length, video frame height, width and number of channels, respectively.
+The goal is to train a detector $\mathcal{F}$ on a subdataset $D_{j}$ built based on a known generation model $g_{j}$, map the video space to the category space $\mathcal{F}$ :::: $\mathcal{X}\rightarrow\mathcal{Y}$, and generalizes to unseen generation models $G_{k}$, where $k\in[1,m],k\neq j$.
 This can be achieved by minimizing classification error of $\mathcal{F}$ on training data $D^{*}_{i}$:
 $\mathop{\arg\min}\limits_{\theta}\mathbb{E}_{(x,y)\in D^{*}_{i}}\mathcal{L}(f(x),y)$ (1)
-Where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$ We implement the evaluation of detection of cross-category generated videos by ensuring that the training set and the test set do not have videos with the same prompt.
+where $\mathcal{L}$ is a loss function, and $\theta$ are the trainable parameters of detector $f$, $D^{*}_{i}$ refers to the training set of $D_{i}$.
+We evaluate the detection of cross-category generated videos by ensuring that the training set and the test set do not contain videos using the same prompt.
 4 Dataset Construction
-To benchmark generated video detectors, we created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ dataset.
-In this section, we mainly introduce the construction of the data set and conduct a comprehensive evaluation of the dataset.
+To benchmark generated video detectors, we have created the first dataset for the generated video detection task: the $GeneratedVideoForensics$ (GVF) dataset.
 4.1 Dataset Composition
-The $GeneratedVideoForensics$ (GVF) dataset $\mathcal{D}$ consists of 964 triples and can be represented as $\mathcal{D}=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $F$ represents multiple fake videos generated by different generation models based on the text prompt $P$.
-In order to examine the generalization of the detectors, we selected four open source generate models with different generation methods and model frameworks: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a), ZeroScope (Sterling., June 2023) and SHOW-1 (Zhang et al., 2023) as generation models for negative samples.
-Therefore, the GVF dataset could be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show1.
-Figure  3 shows an example of one of the triples.
-Figure 4: Data distribution over categories under the “major content (upper) and “attribute control” (lower) aspects.
+The GVF dataset consists of 964 triples and can be represented as $\mathcal{D}=\{(r,p,F)^{(i)}\}_{i=1}^{964}$, where $r$ and $p$ represent the real video and the corresponding text prompt, and $F$ represents the videos generated by the generation models using the text prompt $p$.
+In order to examine the generalizability of the detectors, we have selected four open source generation models employing different generation methods and model frameworks, namely Text2Video-zero [22], ModelScopeT2V [46], ZeroScope [42] and Show-1 [51] as generation models for negative samples.
+Accordingly, the GVF dataset could be divided into four sub-datasets.
+Figure 3 shows an example triple.
+Figure 4: Data distribution across categories under the “major content (upper) and “attribute control" (lower) aspects.
+Figure 5: Performance comparison of the 4 selected models across spatial major contents.
 Prompt.
-To approximate the distribution of videos in real task scenarios, we collected 964 prompts and corresponding real videos from existing open-domain text-to-video datasets: MSVD (Chen & Dolan, 2011) and MSR-VTT (Xu et al., 2016) datasets.
-Following the FTEV (Liu et al., 2023), we consider the collection of prompts from the following four aspects to simulate the sample distribution in real scenarios: the spatial and temporal major content it describes and the spatial and temporal attributes it controls during video generation.
+To approximate the distribution of videos in reality,we have collected 964 prompts and corresponding real videos from open-domain text-to-video datasets: MSVD [6] and MSR-VTT [50].
+Following FTEV [25], we consider the collection of prompts from the following four aspects to simulate the sample distribution in reality: the spatial and temporal content it mainly describes and the spatial and temporal attributes it controls during video generation.
 Spatial content is classified as people, animals, plants, food, vehicles, buildings, and artifacts, with corresponding spatial attributes of quantity, color, and camera view.
-Temporal content and attributes encompass actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order.
-This ensures that our data set is comprehensive and balanced.
-Figure  4 shows the data distribution on categories under different aspects, which shows that the GVF dataset contains enough data to cover different category ranges.
-Figure 5: Manual evaluation results across spatial major contents.
+Temporal content encompasses actions, kinetic motions, fluid motions, and light change, with attributes including motion direction, speed, and event order.
+This ensures that our GVF dataset is comprehensive and balanced.
+Figure 5 illustrates the data distribution across categories under different aspects, which indicates that the GVF dataset contains enough data to cover different category ranges.
 4.2 Dataset Evaluation
-Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation ideas and model architectures, to restore better the challenges posed by invisible generation models.
-Figure 5 shows the performance of videos generated by the above four generation models on different attributes under the FETV  (Liu et al., 2023) standard.
-In order to generate different kinds of content, it is obvious to note that each of these models has its advantages and disadvantages.
-Text2Video-zero, for example, has the best impact on the spatial quality of the video but the worst impact on the temporal continuity.
-On the contrary, ZeroScope has a better effect on temporal continuity but has the worst effect on spatial quality.
-The choice of these video generation models, each possessing distinct advantages and disadvantages, ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
-5 Methodology
-5.1 Probe experiments of STCNN models
-Initially, we considered using a spatiotemporal convolutional model (STCNN) to train the generally generated video detector.
-However, we discovered that these models are only effective on the specific generative models they were trained on and needed more generalizability, failing to discern unseen T2V generation models.
-To analyze what information the STCNN network was learning during discrimination, or rather, which features within the video it was utilizing to distinguish between authentic and fabricated content, we designed a sophisticated probe experiment with two novel datasets under the same model framework: The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to see if the model can still differentiate based solely on the temporal discontinuities of the videos.
-The second dataset was created by randomly selecting a frame from the current test set and replicating it multiple times to see if the model could still differentiate based on the spatial artifacts of the video frame alone.
-The experimental results in table 1 led us to the following conclusion: We found that the model’s attention to spatial artifacts and temporal artifacts was distinct, and for developing a generalized discriminator, they did not play the same role.
-Therefore, we attempt to build detectors based on spatial artifacts and temporal artifacts separately, individually verifying the role each played in the learning process of the universally generated video detector.
-Table 1: Spatiotemporal Convolutional Neural Networks (3DCNN) rely on easily detectable spatial artifacts.
+Among many excellent T2V generation models, we have selected four of the most representative open-source models, including different generation methods and model architectures, to simulate the challenges posed by unseen generation models.
+Figure 5 shows the generation ability of the selected generation models on different attributes under the FETV [25] standard.
+Clearly, each model has its advantages and disadvantages.
+Text2Video-zero, for example, has the best performance on the spatial quality of the video, but the worst performance on the temporal continuity.
+On the contrary, ZeroScope has a better performance on temporal continuity but the worst performance on spatial quality.
+The choice of these video generation models ensures that our dataset can effectively measure the performance of detectors when facing unseen T2V generation models.
+Table 1: Spatiotemporal Convolutional Neural Networks (STCNN) rely on spatial artifacts.
 Method Artifacts Difference
@@
 Mvit 50.48 55.29 72.16 92.59 21.68 37.30
-5.2 Shortcomings of learning from spatial artifacts
-To analyze the role of spatial artifacts, we first work with existing T2I models which have been used as image-level detectors for generated videos, and we test them on the GVF dataset.
+5 Methodology
+5.1 Probe experiments of STCNN models
+Initially, we considered using spatiotemporal convolutional model(STCNN) [5, 12, 11, 10], which are widely used in video classification tasks, to train the generalizable generated video detector.
+However, we discovered that these models are only effective on the specific generation models they were trained on, failing to discern unseen T2V generation models.
+To analyze what information the STCNN network was learning during the training process, or rather, which features within the video it utilized to distinguish between authentic and fabricated content, we designed a probing experiment with two novel datasets under the same model framework.
+The first dataset involves scrambling the frames of all real videos in the current test set, destroying the temporal continuity of the videos, to observe if the model can differentiate based solely on the temporal discontinuities of the videos.
+The second dataset is created by randomly selecting a frame of fake videos from the current test set and replicating it multiple times to observe if the model could differentiate based on the spatial artifacts of the video frame alone.
+The experiment results in Table 1 show that STCNN overly relies on spatial artifacts, degrading the generated video detection into a 2D problem.
+However, the fact STCNN relies on spatial artifacts still does not explain its lack of generalizability to unseen generation models.
+To further explore the possibility of detecting generated video based on spatial artifacts, we turned to the SoTA generated image detection works.
+5.2 Capturing Universal Spatial Artifacts is Difficult
+Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models.
+We transform the current detection of generated images into image level detectors for detecting generated videos and test them on the GVF dataset.
 As shown in Table 2, these detectors demonstrate superior capabilities in capturing the spatial artifacts of specific generation models.
-However, they lack generalization to unseen generation models.
-To find out why these detectors lack generalization to unseen generate models, we visualized the feature space of the detector (Wang et al., 2020) and examined the distribution of four sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, consisting of fake videos from the sub-dataset Text2Video-zero, (iii) Show1, composed of fake videos from the sub-dataset Show-1, and (iv) MP, comprised of fake videos from the sub-dataset ModelScope, (v) ZP, comprised of fake videos from the sub-dataset ZeroScope.
-For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE (Van der Maaten & Hinton, 2008).
-As shown in Figure 6, on the one hand, the detector has ideally classified the real video and the fake video generated by the specific generate model, which means that the detector has fully captured the spatial artifacts of the specific generate model Text2Video-zero  (Khachatryan et al., 2023).
-On the other hand, the detector cannot distinguish fake videos generated by unseen models from real videos.
-Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting other video generation models?
-We illustrate this by visualizing the frequency spectrum of video frames generated by various video generate models.
-As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent; capturing more general spatial artifacts appears to be extremely difficult.
-Figure 6: t-SNE visualization of real and fake video frames associated with four video generation models.
-The feature space used is one trained to distinguish fake videos (Text2Video-zero) from real videos.
-Refer to caption Real
-Refer to caption T2V
-Refer to caption Show1
-Refer to caption MP
-Refer to caption ZP
+However, they also lack generalizability to unseen generation models.
+To find out why these detectors lack generalizability to unseen generate models, we visualized the feature space of the detector [47] and examined the distribution of five sub-datasets: (i) Real, comprising real videos from the test set, (ii) T2V, comprising fake videos from the sub-dataset Text2Video-zero, (iii) Show1, comprising fake videos from the sub-dataset Show-1, and (iv) MP, comprising fake videos from the sub-dataset ModelScope, (v) ZP, comprising fake videos from the sub-dataset ZeroScope.
+For each video, we utilized the detector trained on (i) and (ii) to extract their feature representations corresponding to each frame and subsequently visualized these through t-SNE [44].
+Real
+T2V
+Show1
+MP
+ZP
 Figure 7: Average spectrum of video frames for real videos and fake videos generated by different video generation models.
+As shown in Figure 6, on the one hand, the detector has successfully classified the real video and the fake video generated by the seen generate model (Text2Video-zero), indicating that the detector has fully captured the spatial artifacts of the videos generated by the seen generate model.
+On the other hand, the detector cannot distinguish fake videos generated by unseen models from the real videos.
+Why are the spatial artifacts captured by the detector in (ii) not helpful for detecting videos generated by other video generation models?
+We illustrate this by visualizing the frequency spectrum of video frames generated by the models.
+As shown in Figure 7, although the generated videos all have spatial artifacts, such as parallel slashes appearing in T2V and bright spots appearing in Show-1, these spatial artifacts are not consistent.
+Therefore, capturing universal spatial artifacts appears to be extremely difficult.
 Figure 8: Overview of the DeCoF framework.
-We first get real video and generated video features by the pre-trained model CLIP:VIT, which is designed to eliminate the impact of spatial artifacts on capture temporal artifacts, and then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of real and fake video.
-5.3 Eliminate impact of spatial artifacts
-There is no doubt that real video frames are continuous, and we can even predict any frame of the video through other frames.
-However, mismatches are prone to occur between the generated video frames, resulting in the inability to meet the continuity of the video frames, that is, temporal artifacts of the generated video.
-Given a set of video frames $V=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a particular function that measures the continuity between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video.
-After analyzing the results of the above exploration, it appears that the detection model tends to learn spatial artifacts that are easy to capture but do not have generalization properties.
-The difficulty lies in how to eliminate the impact of spatial artifacts on capturing temporal artifacts.
-Intuitively, if we can find a particular mapping function $g$ with the following properties: The distance between frame mappings is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts.
-In fact, we choose ViT-L/14  (Dosovitskiy et al., 2020) trained for the task of image-language alignment as mapping function $g$, then by mapping the video frames $F_{i}$ to $g(F_{i})$, the decoupling of temporal artifacts and spatial artifacts can be achieved.
-As a result, the original problem can be expressed as $f(g(F_{1}),g(F_{2}),\cdots,g(F_{L}))=y$.
+We first get real video and generated video features using the pre-trained CLIP:VIT, to eliminate the impact of spatial artifacts on capturing temporal artifacts.
+Then a verification module consisting of two transformer layers and one MLP head is used to learn the differences between frame consistency of the real and fake videos.
+5.3 Detection based on Temporal Artifacts
+The recent video generation models convert the T2I model into the T2V model through additional constraints, which have common drawbacks in controlling video continuity in two aspects: 1) the normativity of motion, not only do visible movements violate to physical laws, but more importantly, there are irregular movements of invisible detail pixels, 2) there is a slight pixel shift in the relatively static local background.
+Therefore, we propose to develop a detection method based on temporal artifacts.
+Given a set of video frames $F=\{F_{1},F_{2},\cdots,F_{L}\}$, we define the method to determine the authenticity of a video based on its temporal artifacts as $f(F_{1},F_{2},\cdots,F_{L})=y$, where $f$ is a specific function that measures the consistency between video frames, $y\in Y=\{0,1\}$ is the authenticity label of the video.
+It appears that the detection model tends to learn spatial artifacts (Section 5.1).
+Therefore, we propose to eliminate the impact of spatial artifacts in order to help the model focus on capturing temporal artifacts.
+Intuitively, we want to find a particular mapping function $M$ with the following properties: the distance between mapping frames is negatively related to the image similarity, and the mapping function is insensitive to spatial artifacts.
+Hence, we choose ViT-L/14 [8] trained for the task of image-language alignment as mapping function $M$.
+Then by mapping the video frames $F_{i}$ to $M(F_{i})$, We have implemented decoupling the temporal and spatial artifacts because the mapping function is not trained to capture spatial artifacts, but only encodes the image into a feature space with excellent properties, where the distance between features is inversely proportional to the image similarity.
+As a result, the original problem can be expressed as $f(M(F_{1}),M(F_{2}),\cdots,M(F_{L}))=y$.
 Model framework.
-In order to learn video frame consistency for capturing temporal artifacts, we design our detection with the Consistency of video Frames (DeCoF), in which the critical verification module consists of two transformer layers and an MLP head, As shown in Figure  8.
-More specifically, given a video input $V$, The frozen parameter image encoder (ViT-L/14) will extract the features of the video frames and generate a feature sequence $S=\{g(F_{1}),g(F_{2}),\cdots,g(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature.
+In order to learn video frame consistency for capturing temporal artifacts, we design a detection method based on the Consistency of video Frames (DeCoF), in which the core verification module consists of two transformer layers and an MLP head as shown in Figure 8.
+More specifically, given a video input $V$, the frozen parameter image encoder (ViT-L/14) extracts the features of the video frames and generates a feature sequence $S=\{M(F_{1}),M(F_{2}),\cdots,M(F_{L})\}$ with a shape of $L*D$, where $D$ represents the dimension of each feature.
 Next, the position encoding and a learnable class embedding representing the classification output are prepended to the sequence $S$ and then passed through a sequential module composed of two layers of transformers.
-Finally, a classification header is appended to the transformer layer, resulting in a probability distribution over the target class, and we train our model via a simple cross-entropy loss.
+Finally, a classification header is appended to the transformer layer, outputting a probability distribution over the target class.
+We train our model via a simple cross-entropy loss.
 In addition to decoupling temporal and spatial artifacts, our approach significantly reduces computational complexity and memory requirements since we only need to learn anomalies between video frame features.
-Overall, our approach is simple yet effective.
+As such, we achieve a simple yet effective approach.
 6 Experiments
-In this section, we introduce the experiment setup and provide extensive experimental results to demonstrate the superiority of our approach.
-Table 2: Comprehensive comparison of our method with previous detectors and spatiotemporal convolutional networks, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model.
-* Denotes our retraining on a subset of generated video forensics with official code.
+6.1 Experimental Setup
+Data pre-processing.
+We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalizability of the detector.
+We take first 8 frames of each video and center-crop each frame according to its shortest side, then resize it to 224 × 224.
+During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur are also used.
+Evaluation metrics.
+Following previous methods [47, 48, 13], we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors.
+The threshold for computing accuracy is set to 0.5 following [47, 48].
+For image-based methods, we average frame-level predictions into corresponding video-level predictions.
+Baseline.
+(i) CNNDet [47] detects images generated by GAN.
+Through a large amount of data enhancement, the model has generalizability on images generated by various GANs.
+(ii) DIF [39] is a few-shot learning method for image synthesis.
+(iii) DIRE [48] proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models.
+(iv) Lgard [43] uses a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts.
+(v) I3D [5] introduces a new Inflated 3D ConvNet that is based on 2D ConvNet inflation.
+(vi) Slow [12] proposes a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames.
+(vii) X3D [11] designs a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy.
+(viii) Mvit [10] combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition.
+(ix) UFDetect(UniversalFakeDetect) [29] detects generated images through pretrained CLIP:ViT visual encoders and nearest neighbors.
+(x) EVR [24] uses the CLIP backbone with frozen parameters and multi layer spatiotemporal Transformer decoder for video action recognition.
+6.2 Evaluation
+We conduct comprehensive evaluations to answer the following questions:
+Table 2: Comprehensive comparison of our method with previous detectors and STCNN, including CNNDet, DIF, DIRE, Lgard, I3D, Slow, X3D and Mvit, were tested on our GVF dataset using the officially provided model.
+* Denotes we retrained it on a subdataset of GVF dataset with official code.
 We report ACC (%) and AP (%) (ACC and AP in the Table).
-Method Training dataset Generate mode Detection Level Tested generate video models Total Avg.
+Method Training dataset Generation mode Detection Level Video generation models Total Avg.
 Text2Video-zero Show1 ModelScope ZeroScope
 ACC AP ACC AP ACC AP ACC AP ACC AP
-CNNDet(0.5) LSUN ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
-CNNDet(0.1) LSUN ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
-DIF SDV2 Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
-DIRE LSUN-B Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
-Lgard LSUN ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
+CNNDet(0.5) - ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
+CNNDet(0.1) - ProGAN Image 49.93 46.38 50.06 46.72 50.19 42.97 50.06 55.27 50.06 47.84
+DIF - Diffusion Image 35.10 47.96 45.90 53.42 31.40 46.91 27.50 46.59 34.98 48.72
+DIRE - Diffusion Image 63.14 63.45 51.80 51.70 51.48 51.03 46.71 42.96 53.28 52.29
+Lgard - ProGAN Image 53.09 64.78 59.70 68.42 40.21 38.87 47.94 51.80 50.24 55.97
+UFDetect - ProGAN Image 50.13 63.99 53.09 74.42 52.13 75.55 58.89 84.27 53.56 74.56
 CNNDet (0.5)* Text2Video-zero T2V Image 98.64 99.97 51.93 61.64 51.54 68.91 54.96 72.51 64.27 75.76
 CNNDet (0.1)* Text2Video-zero T2V Image 96.84 99.99 53.93 60.99 55.34 68.87 60.69 73.80 66.70 75.91
+DIRE* Text2Video-zero T2V Image 99.17 99.99 49.93 56.74 52.83 77.58 51.28 67.12 63.30 75.36
 Lgard* Text2Video-zero T2V Image 98.45 100.00 56.19 78.91 51.03 66.30 54.64 75.93 65.08 80.29
-DIRE* Text2Video-zero T2V Image 99.17 99.99 49.93 56.74 52.83 77.58 51.28 67.12 63.30 75.36
 I3D* Text2Video-zero T2V Video 98.00 99.87 49.75 70.41 49.25 71.43 50.50 51.85 61.88 73.39
@@
 Mvit* Text2Video-zero T2V Video 98.42 99.97 49.96 45.93 49.49 38.98 51.68 32.60 62.39 54.37
+EVR* Text3Video-zero T2V Video 97.50 99.93 78.5 90.96 59.75 77.97 64.75 85.46 67.67 88.58
 DeCoF Text2Video-zero T2V Video 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
@@
 DeCoF ZeroScope T2V Video 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
-6.1 Experimental Setup
-Data pre-processing.We only train on a certain sub-dataset of GVF and test on the remaining sub-datasets to evaluate the generalization of the detector.
-We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224.
-During the training process, some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
-Evaluation metrics.
-Following previous methods  (Wang et al., 2020, 2023b; Frank et al., 2020), we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors.
-The threshold for computing accuracy is set to 0.5 following  (Wang et al., 2020, 2023b).
-For image-based methods, we average frame-level predictions into corresponding video-level predictions.
-Baseline.
-(i) CNNDet (Wang et al., 2020) proposes a model for detecting images generated by CNN.
-Through a large amount of data enhancement, the model has generalization on various images generated by CNN.
-(ii) DIF  (Sinitsa & Fried, 2023) proposes a few-shot learning method for image synthesis.
-(iii) DIRE (Wang et al., 2023b) proposes a new image representation method called Diffusion Reconstruction Error (DIRE), which measures the error between the input image and its reconstruction via a pre-trained diffusion model to detect images generated by diffusion models.
-(iv) Lgard (Tan et al., 2023) used a pre-trained CNN model as a transformation model to convert images into gradients and used these gradients to present generalized artifacts.
-(v) I3D (Carreira & Zisserman, 2017) introduce a new Inflated 3D ConvNet that is based on 2D ConvNet inflation.
-(vi) Slow (Feichtenhofer et al., 2019) proposed a dual-path model, but we only adopted the Slow Pathway, which is designed to capture the semantic information provided by a small number of sparse frames.
-(vii) X3D (Feichtenhofer, 2020) designed a smaller and lighter structure and expanded it in different dimensions to achieve a trade-off between efficiency and accuracy.
-(viii) Mvit  (Fan et al., 2021) combines the basic idea of multiscale feature hierarchies with transformer models for video and image recognition.
-6.2 Detailed Evaluation
-We conduct comprehensive and challenging experiments to answer the following questions:
+Q1. How does our detector generalize to unseen T2V models compared to previous detectors?
+Table 2 compares CNNDet [47], DIF [39], DIRE [48], Lgard [43], UFDetect [29], using the pre-trained weights provided by the official repository.
+Although these detectors have good performance in detecting generated images, their performance drops significantly on the task of detect generated videos, with both ACC and AP below 60%.
+As a result of retraining the above five detectors on the sub-dataset Text2Video-zero, their performance on seen video generation models greatly improved, but they still performed poorly for unseen generation models, with ACC and AP scores still lower than 60%, and the retrained STCNN presents similar results.
+On the contrary, our DeCoF has excellent generalizability for various unseen generation models.
+Compared with previous video classification methods based on pretrained CLIP [24], our method has still made significant progress.
+Specifically, on the task of generated video detection, DeCoF outperforms previous detectors by 13.77% $\sim$ 31.38% in ACC and 8.44% $\sim$ 49.18% in AP.
+We also conducted cross-dataset evaluations.
+As shown in Table 2, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%.
 Figure 9: Robustness to unseen perturbations.
 The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression.
-We test the detectors trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
-Q1. How does our detector perform at detecting unseen T2V models compared to previous detectors?
-As mentioned in Section 5.2, we found that previous detectors exhibit extraordinary capabilities when detecting specific generative models but lack generalization when unseen generative models are present.
-In the table 2, we quantitatively evaluate CNNDet (Wang et al., 2020), DIF (Sinitsa & Fried, 2023), DIRE (Wang et al., 2023b), Lgard (Tan et al., 2023), using the pre-trained weights provided by the official repository.
-Although these detectors have surprising performance in detecting generated images, their performance drops significantly on the task of detecting generated videos, with both ACC and AP below 60%.
-As a result of retraining the above four detectors on the sub-dataset Text2Video-zero, their performance on specific video generation models greatly improved, but they still performed poorly for unseen generation models, and ACC and AP scores are still lower than 60%, and even the retrained spatiotemporal convolutional network displays similar results.
-On the contrary, compared with previous detectors, our generated video detection method has excellent generalization for various unseen generation models.
-Specifically, on the task of generating video detection, the ACC increased by 17.74% $\sim$ 31.38% and the AP by 16.73% $\sim$ 49.18%.
-We also conducted cross-dataset evaluations.
-Table  2 shows that on the task of generating video detection, the ACC of the model trained on the other three data sets is much higher than that of the model trained on Text2Video-zero by 5.25% $\sim$ 13.44%.
-This is because Text2Video-zero has the worst temporal continuity in all four T2V generation models.
-Table 3: Performance of DeCoF on commercial generation models.
-Training dataset Tested generate video models Total Avg.
-Gen-2 Pika
-ACC AP ACC AP ACC AP
-Text2Video-zero 99.49 100.00 99.49 99.98 99.49 99.99
-Show1 98.46 99.96 98.46 99.85 98.46 99.91
-ModelScope 98.97 99.99 96.89 99.73 97.93 99.86
-ZeroScope 98.97 99.99 94.84 99.52 96.91 99.76
+We test the detector trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
 Q2. How does our detector maintain robustness against noise and disturbances?
@@
 The perturbations are added under three levels for Gaussian blur ($\sigma$ = 1, 2, 3) and five levels for JPEG compression (quality = 90, 80, 70, 60, 50).
-Figure  9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations.
-Although the model’s performance in the face of perturbations is inconsistent due to differences in training data sets, Our method remains effective under any degree of arbitrary perturbation, and the AP of models trained on any sub-dataset is greater than 80%.
+Figure 9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations.
+Although experiencing some performance inconsistencies due to differences between training datasets, our method remains effective under all degrees of arbitrary perturbation, with the AP higher than 80%.
 Q3. How does our detector perform on commercial generation models?
-To further evaluate the effectiveness of our detection method, we constructed two additional test datasets using two well-known commercial proprietary T2V generation models, Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023).
-Based on the datasets generated by the four open-source T2V models shown in Table  3, the detector was trained on Text2Video-zero, Show1, ModelScope, and ZeroScope.
-When confronted with the commercial generation model, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2, demonstrating the strong generalization capability of our model DeCoF for unseen T2V models.
-Even when trained on specific T2V models, the model is highly accurate in detecting newly generated video.
+Table 3: Performance of DeCoF on commercial generation models.
+Training dataset Video generation models
+Gen-2 Pika Sora Veo
+ACC AP ACC AP ACC AP ACC AP
+Text2Video-zero 99.49 100.00 99.49 99.98 96.87 99.70 96.43 100.00
+Show1 98.46 99.96 98.46 99.85 94.79 98.26 92.86 100.00
+ModelScope 98.97 99.99 96.89 99.73 94.79 98.83 89.29 98.47
+ZeroScope 98.97 99.99 94.84 99.52 89.58 97.92 67.86 94.90
+To further evaluate the effectiveness of our detection method, we constructed four additional test datasets using well-known commercial proprietary T2V models, Pika [32], Gen-2 [15], Sora [41] and Veo [45].
+As shown in Table 3, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2.
+Due to regional restrictions, we did not qualify for testing Sora and Veo directly, but we constructed test sets for 48 public videos released by Sora and 14 public videos released by Veo, respectively.
+In this test, we achieved no less than 89.58% ACC and no less than 97.92% AP on Sora, no less than 67.86% ACC and no less than 94.90% AP on Veo, demonstrating the strong generalizability of our model DeCoF for unseen commercial T2V models.
 7 Conclusions and Future Work
-The paper presents a groundbreaking investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content.
-We construct the first public dataset for real and fake video detection, providing a solid benchmark for community research.
-At the same time, we propose a simple yet effective detection model, which has proven its strong generalizability and robustness on multiple open-source and commercial closed-source models.
+The paper presents the first investigation on generated video detection, with the goal of tackling the societal concern of validating the authenticity of video content.
+We construct the first public dataset for video detection, providing a solid benchmark for community research.
+At the same time, we propose a simple yet effective detection model, and prove its strong generalizability and robustness on multiple open-source and commercial closed-source models.
 We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
-8 Impact Statements
-Our research focuses on using Machine Learning to detect generated videos.
-We developed the DeCoF model and shared our dataset to support detecting synthetic videos.
-These works are crucial to protect digital content and prevent the spread of disinformation.
-Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology.
-We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity.
-We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
-Appendix A $GeneratedVideoForensics$ Dataset
-A.1 Dataset Composition
-The following section provides a more detailed description of how the generated video forensics dataset was constructed and evaluated. The entire GVF dataset is generated in two steps:
-1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  (Chen & Dolan, 2011) and MSR-VTT  (Xu et al., 2016) datasets.
-When selecting samples, we followed the work of FETV  (Liu et al., 2023), Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios.
-More specifically, in the work of FETV, they classify each prompt controlling video generation from two aspects: the main content it describes and the attributes it controls during video generation.
+Appendix A Appendix
+This supplementary material provides:
+  • Sec. B: A detailed description of dataset construction, including dataset structure diagrams, as well as specific examples of prompts and videos.
+  • Sec. C : Various parameter details for model training, hyperparameter settings, etc.
+  • Sec. D : The impact of selecting different mapping functions.
+  • Sec. E : The impact of video frame count.
+  • Sec. F : The impact of video frame sampling interval.
+  • Sec. G : More details on Sora.
+  • Sec. H : Spatial artifacts vs Temporal artifacts.
+  • Sec. I : Impact Statements.
+  • Sec. J : Limitation.
+  • Sec. K : More visual examples.
+Appendix B $GeneratedVideoForensics$ Dataset
+B.1 Dataset Composition
+The following section provides a more detailed description of how the $GeneratedVideoForensics$ was constructed and evaluated. The entire GVF dataset is generated in two steps:
+1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  [6] and MSR-VTT  [50] datasets.
+When selecting samples, we followed the work of FETV  [25], Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios.
+More specifically, in the work of FETV, they classify each prompt from two aspects: the main content it describes and the attributes it controls during video generation.
 The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations.
@@
 Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
-2) Selection of generation models: Although there are many high-quality generative models on the market, we do not want to choose the best.
-On the contrary, we have selected four generative models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of Generative models appear, how do we deal with the problems caused by invisible generative models?
-In the end, the four most representative generative models were selected as our generative models: Text2Video-zero (Khachatryan et al., 2023), ModelScopeT2V (Wang et al., 2023a),ZeroScope(Sterling., June 2023),Show1 (Zhang et al., 2023).
-At the same time, we also selected the two most popular models as the secret test dataset: Pika (Pika, December 2023) and Gen-2 (Gen-2, December 2023).
+2) Selection of generation models: Although there are many high-quality generation models on the market, we do not want to choose the best.
+On the contrary, we have selected four generation models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of generation models appear, how do we deal with the problems caused by unseen generation models?
+In the end, the four most representative generation models were selected as our generation models: Text2Video-zero [22], ModelScopeT2V [46], ZeroScope[42], Show1 [51].
+At the same time, we also selected the three most popular models as the secret test dataset: Pika [32], Gen-2 [15], Sora[41] and Veo [45].
+For Sora and Veo, we provide detailed information in the section G
 If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$.
@@
 Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
-A.2 Dataset Evaluation
-As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV (Liu et al., 2023).
+B.2 Dataset Evaluation
+As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV [25].
 Table 4: Some examples of prompts.
@@
 Fluid Motions Speed a bucket of slime is quickly poured on top of a woman’s head
-Figure 10: Data distribution over categories under the “major content” (upper) and “attribute control” (lower) aspects.
 Attribute statistics of prompts.
@@
 At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different.
-You may ask why time main content and time attributes are so rare.
-In fact, not every prompt in the MSVD and MSR-VTT datasets has a time description, and we only ensure the existence of this type of prompt in the data set.
 Here, we also provide some specific examples of prompts, As shown in the table 4.
+Figure 10: Data distribution over categories under the “major content” (upper) and “attribute control” (lower) aspects.
 Evaluation of different generation models.
-Here, we present our assessment of four generation models according to the FETV standard (Liu et al., 2023).
+Here, we present our assessment of four generation models according to the FETV standard [25].
 We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames.
@@
 Therefore, it is more accurate to assess controllability using fine alignment when examining a specific attribute, whereas global alignment is a reasonable approximation.
-Appendix B Parameter Description
+Figure 11: The manual evaluation provided results across major static and temporal contents and attributes to control video-text alignment.
+Appendix C Parameter Description
 In this section, we provide a detailed description of various parameters of the experiment, including the hyperparameters of the model and the details of its training.
@@
 During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets.
-The numbers of the training set, validation set, and test set are 771, 97, and 96, respectively.
-We freeze the parameters of CLIP:ViT (Dosovitskiy et al., 2020), making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9.
+The numbers of the training set, validation set, and test set are 771, 96, and 97, respectively.
+We freeze the parameters of CLIP:ViT [8], making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9.
+All experiments were completed on two RTX 3090 graphics cards.
 Data pre-processing.
 All the following experiments are performed on our GVF dataset.
-In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 (Zhang et al., 2023), ModelScope (Wang et al., 2023a) and ZeroScope (Sterling., June 2023) Since the resolution and length of the videos generated by the four generative models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224.
+In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 [51], ModelScope [46] and ZeroScope [42] Since the resolution and length of the videos generated by the four generation models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224.
 During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224.
@@
 Our model consists of two layers of transformer and an MLP head is an adaptation of the sequence to sequence trans- former, where the dimension and feature dimension of MLP is 768, path_size is 32, the dropout parameter is 0.1, and the number of attention heads is 4.
-Table 5: A comparison of the impact of different mapping functions.
-Mapping Function Pretrain Tasks Tested generate video models Total
-Avg.
-Text2Video-zero Show1 ModelScope ZeroScopre
-ACC AP ACC AP ACC AP ACC AP ACC AP
-ViT-L/14 CLIP 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
-Resnet50 CLIP 96.92 99.82 52.60 80.66 57.22 80.19 50.00 67.65 64.19 82.08
-ViT-B/16 ImageNet
-Resnet50 ImageNet
-Figure 11: The manual evaluation provided results across major static and temporal contents and attributes to control video-text alignment.
-Appendix C Different Mapping Functions
+Appendix D Different Mapping Functions
 In this section, we examine the impact of different mapping functions.
@@
 First, we discussed the two different tasks of ImageNet and CLIP.
-Interestingly, the encoder trained on ImageNet caused this problem not to converge.
+Compared to ImageNet’s image classification task, CLIP’s image encoder trained on text image task alignment is a better mapping function.
 We considered two reasons: one is the size of the training data, and the other is due to the difference in the task itself; We do not want the model to classify images into specific items or biological categories but rather cluster similar images.
-At the same time, we compared the impact of different model frameworks.
-Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function.
+At the same time, we compared the impact of different model frameworks: Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function.
 Overall, the mapping function largely determines the effectiveness of our approach.
-Appendix D More Visual Examples
+Table 5: A comparison of the impact of different mapping functions.
+Mapping Function Pretrain Tasks Video generation models Total
+Avg.
+Text2Video-zero Show1 ModelScope ZeroScope
+ACC AP ACC AP ACC AP ACC AP ACC AP
+ViT-L/14 CLIP 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
+Resnet50 CLIP 96.92 99.82 52.60 80.66 57.22 80.19 50.00 67.65 64.19 82.08
+ViT-B/16 ImageNet 98.00 99.79 65.39 87.65 54.70 76.39 55.57 76.1 68.41 84.98
+Resnet50 ImageNet 99.5 99.96 54.2 81.9 51.43 67.97 49.93 58.65 63.76 77.12
+Appendix E The Impact of Video Frame Number
+Here we discuss the impact of the number of video frames used for detection.
+Table 6: A comparison of the impact of different frame numbers.
+Training dataset Frame Numbers Video generation models Total
+Avg.
+Text2Video-zero Show1 ModelScope ZeroScope
+ACC AP ACC AP ACC AP ACC AP ACC AP
+Text2Video-zero 8 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
+Show1 8 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
+ModelScope 8 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
+ZeroScope 8 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
+Text2Video-zero 16 99.50 99.99 60.70 95.13 65.55 93.55 63.27 95.38 72.25 96.01
+Show1 16 97.86 99.69 95.93 99.04 93.73 98.13 94.30 98.88 95.45 98.93
+ModelScope 16 100.00 100.00 81.25 95.41 98.36 99.55 95.36 99.19 93.74 98.53
+ZeroScope 16 97.43 99.83 80.75 96.10 82.18 96.26 96.93 99.72 89.32 97.97
+Text2Video-zero 24 99.50 100.00 57.57 95.14 55.7 94.10 60.18 96.32 68.23 96.39
+Show1 24 97.80 99.98 94.36 99.10 89.66 97.49 93.73 98.93 93.88 98.87
+ModelScope 24 100.00 100.00 74.68 95.57 97.00 99.32 90.73 99.30 90.60 98.54
+ZeroScope 24 98.43 99.96 83.32 97.00 85.02 96.56 97.50 99.78 91.06 98.32
+In our work, we selected 8 frames for generated video detection for two reasons: on the one hand, the current generation model still has this deficiency in generating long videos.
+Even the videos currently displayed by Sora can only generate about 300-600 frames, so we believe that 8 frames is an acceptable choice.
+On the other hand, in order to highlight the effectiveness of our method, we chose a challenging number of 8 frames, and a large number of experiments have also proven that our method indeed performs very well.
+We conducted experiments on the impact of frames 8, 16, and 24 on detection, as shown in the table 6.
+We found that the performance of 16 frames is generally better than that of 8 frames, except for the models trained on the Text2Video zero subdataset, which is consistent with our previous evaluation results of several generation models.
+Text2Video zero has the worst time continuity.
+In addition, the effect will become worse at 24 frames.
+One possible reason is that as the sequence increases, the temporal artifacts of the generated video become more unique, which weakens the model’s generalizability.
+Appendix F The Impact of Video Frame Sampling Interval.
+In this section, we discussed the impact of sampling interval on detection results. In previous experiments, we usually took the first 8 frames of the video for detection.
+Unlike previous experiments, this experiment was conducted on the latest generation model Sora, as Sora has made significant progress in generating frames compared to other generation models, making it easier for us to comprehensively discuss the impact of sampling strategies.
+As shown in the table 7, facing three different sampling strategies: 1) Take the first 8 frames; 2) Take 8 frames evenly from all frames; 3) Take 8 frames evenly from the first 300 frames; Our method did not show a significant decrease, which we attribute to two reasons: 1) The prompts we collected were comprehensive enough to cover the speed of video changes well.
+2) The duration of the test video is still around 10 seconds, so the impact of sampling interval is not very significant.
+For longer generated videos, the performance degradation caused by changes in sampling intervals will be more severe.
+Table 7: The impact of different sampling strategies
+Training dataset Sampling strategy Sora
+ACC AP
+Text2Video-zero First 8 frames 96.87 99.70
+Show1 First 8 frames 94.79 98.26
+ModelScope First 8 frames 94.79 98.83
+ZeroScope First 8 frames 89.58 97.92
+Text2Video-zero Take 8 frames evenly from all frames 96.87 99.48
+Show1 Take 8 frames evenly from all frames 94.79 98.61
+ModelScope Take 8 frames evenly from all frames 94.79 98.96
+ZeroScope Take 8 frames evenly from all frames 88.54 98.31
+Text2Video-zero Take 8 frames evenly from first 300 frames 97.91 99.65
+Show1 Take 8 frames evenly from first 300 frames 94.79 98.48
+ModelScope Take 8 frames evenly from first 300 frames 93.75 98.74
+ZeroScope Take 8 frames evenly from first 300 frames 90.62 98.22
+Appendix G More Details on Sora.
+G.1 Test data on Sora
+Due to regional restrictions, we were not eligible to test Sora.
+But we tested all of the 48 videos officially displayed by Sora.
+The test dataset consists of 48 official videos released by Sora and 48 real videos from the original test set, the test on Veo are similar.
+G.2 The detection performance of human testers on Sora
+We invited three testers to evaluate the authenticity of 96 videos (48 real and 48 Sora generated) from the previously mentioned Sora test set. The experimental results are shown in the table 8.
+Table 8: Compared to the test results of human testers
+Training dataset Sora
+TP+TN ACC
+Tester 1 57(48+9) 59.37
+Tester 2 66(47+19) 68.75
+Tester 3 60(48+12) 62.5
+Text2Video-zero 93 96.87
+Show1 91 94.79
+ModelScope 91 94.79
+ZeroScope 86 89.58
+The results of DeCoF in testing far exceed the level of human judgment.
+At the same time, we also need to point out that interviews with testers after testing revealed that their judgment on the videos generated by Sora was artificially high, as there were 4-5 flashy videos in the officially released videos that did not match the results that could appear in the real world, such as pirate ships in coffee cups.
+Meanwhile, according to the testers, their judgment is based on whether there is irregular movement (such as objects appearing out of thin air) or object threading (such as overlapping basketball and basket) and abnormal distortion (such as human fingers or limbs of moving animals) in the generated video.
+Appendix H Spatial artifacts vs Temporal artifacts.
+H.1 Detect generated videos from temporal artifacts
+Although our method attempts to detect generated videos from temporal artifacts, there is a question: is it possible that CLIP: ViT has differences in the features between the generated video and the real video frames when extracting image features, resulting in better detection performance?
+In other words, is our method really detect generated videos from temporal artifacts?
+To prove this, we repeated the previous probe experiment.
+As shown in the table 9, our model has also performed well in detecting test data without spatial artifacts.
+On the other hand, for the second set of test data in the previous probe experiment, our model has also successfully detected it.
+We believe that repeated stillness in multiple frames is also a type of temporal artifact to some extent.
+Table 9: The proof of our method detect generated videos from temporal artifact
+Method Video generation models
+Temporal Spatial
+ACC AP ACC AP
+I3D 51.17 51.61 86.67 98.61
+Slow 50.04 50.37 93.81 99.51
+X3D 50.04 54.31 78.39 92.84
+Mvit 50.48 55.29 72.16 92.59
+Ours 98.46 99.80 98.46 99.66
+H.2 Spatial artifacts vs Temporal artifacts.
+In this paper, we attempt to detect generated video from the perspective of temporal artifacts, but this does not mean that it is impossible to detect generated video from spatial artifacts.
+In fact, the difficulty of detecting spatial artifacts in generated videos lies in how to extract more general artifacts, as different generation models generate videos with different resolutions, color brightness, contrast, and so on.
+Appendix I Impact Statements
+Our research focuses on detecting generated videos.
+We developed the DeCoF model and shared our dataset to support detecting generated videos.
+These works are crucial to protect digital content and prevent the spread of disinformation.
+Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology.
+We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity.
+We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
+Appendix J Limitation
+In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models.
+However, there are various methods of video forgery, such as Deepfake and malicious editing, in which our methods may experience significant performance degradation or inapplicability.
+Taking Deepfake detection as an example, our method does not add additional detection modules to the face.
+In addition, Deepfake videos are often locally tampered with, and the generated videos are generated as a whole.
+Our work focuses on detecting generated videos, which has a more diverse range of forgery scenarios and more flexible forgery techniques.
+Appendix K More Visual Examples
 Here, we provide more examples of generated videos. Figures 12 to Figures 15 are examples of real videos and generated videos corresponding to prompts.

### v4

@@
-DeCoF: Generated Video Detection via Frame Consistency: The First Benchmark Dataset
+DeCoF: Generated Video Detection via Frame Consistency
 Abstract
@@
 Extensive experiments demonstrate the efficacy of DeCoF in detecting videos generated by unseen video generation models and confirm its powerful generalizability across several commercially proprietary models.
-Our code and dataset will be released at https://anonymous.4open.science/r/DeCoF-8394.
+Our code and dataset will be released at https://github.com/wuwuwuyue/DeCoF.
 1 Introduction
@@
 Figure 15: Real videos and videos generated by four generation models according to prompt: A boy rides a bike on a fence.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We have accurately reflected the contribution and scope of the paper in the abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: As shown in section I of the appendix, we discussed the limitations of the work.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+[…]
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer:[Yes]
+  19. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete experimental configuration and model architecture.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24.
+Justification: In the abstract, we provided the open address of the code and dataset, but the section on commercial models in the dataset is not yet complete.
+For example, due to limitations in testing qualifications, the construction of the complete Sora and Veo test sets has not yet been completed.
+At the same time, the code has not been fully organized and improved, we will gradually open up our project.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete experimental configuration and model architecture
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+[…]
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete sufficient information on the computer resources.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+[…]
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. As shown in section I of the appendix, we discussed the broader impacts of the work.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generation models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+[…]
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+[…]
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We have provided a detailed description of the construction process of the dataset in our paper, and we will then open source the dataset and detection methods.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [No]
+  69. Justification: Our work only includes one evaluation task involving human subjects, mainly to distinguish the authenticity of videos, which is a very simple task.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer:[N/A]
+[…]
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v5

@@
 Figure 15: Real videos and videos generated by four generation models according to prompt: A boy rides a bike on a fence.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We have accurately reflected the contribution and scope of the paper in the abstract and introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: As shown in section I of the appendix, we discussed the limitations of the work.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-[…]
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer:[Yes]
-  19. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete experimental configuration and model architecture.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24.
-Justification: In the abstract, we provided the open address of the code and dataset, but the section on commercial models in the dataset is not yet complete.
-For example, due to limitations in testing qualifications, the construction of the complete Sora and Veo test sets has not yet been completed.
-At the same time, the code has not been fully organized and improved, we will gradually open up our project.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete experimental configuration and model architecture
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-[…]
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: In section  6.1 of the main text and section  C of the appendix, we provide the complete sufficient information on the computer resources.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-[…]
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. As shown in section I of the appendix, we discussed the broader impacts of the work.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generation models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-[…]
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-[…]
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We have provided a detailed description of the construction process of the dataset in our paper, and we will then open source the dataset and detection methods.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [No]
-  69. Justification: Our work only includes one evaluation task involving human subjects, mainly to distinguish the authenticity of videos, which is a very simple task.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer:[N/A]
-[…]
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v6

@@
 Temporal Spatial
-ACC AP ACC AP ACC AP
+ACC AUC ACC AUC ACC AUC
 I3D 51.17 51.61 86.67 98.61 35.50 47.00
@@
 Evaluation metrics.
-Following previous methods [47, 48, 13], we mainly report accuracy (ACC) and average precision (AP) in our experiments to evaluate the detectors.
+Following previous methods [47, 48, 13], we mainly report accuracy (ACC) and area under curve (AUC) in our experiments to evaluate the detectors.
 The threshold for computing accuracy is set to 0.5 following [47, 48].
@@
 * Denotes we retrained it on a subdataset of GVF dataset with official code.
-We report ACC (%) and AP (%) (ACC and AP in the Table).
+We report ACC (%) and AUC (%) (ACC and AUC in the Table).
 Method Training dataset Generation mode Detection Level Video generation models Total Avg.
 Text2Video-zero Show1 ModelScope ZeroScope
-ACC AP ACC AP ACC AP ACC AP ACC AP
+ACC AUC ACC AUC ACC AUC ACC AUC ACC AUC
 CNNDet(0.5) - ProGAN Image 49.61 41.47 49.74 47.58 49.61 43.40 49.94 59.85 49.73 48.08
@@
 Table 2 compares CNNDet [47], DIF [39], DIRE [48], Lgard [43], UFDetect [29], using the pre-trained weights provided by the official repository.
-Although these detectors have good performance in detecting generated images, their performance drops significantly on the task of detect generated videos, with both ACC and AP below 60%.
-As a result of retraining the above five detectors on the sub-dataset Text2Video-zero, their performance on seen video generation models greatly improved, but they still performed poorly for unseen generation models, with ACC and AP scores still lower than 60%, and the retrained STCNN presents similar results.
+Although these detectors have good performance in detecting generated images, their performance drops significantly on the task of detect generated videos, with both ACC and AUC below 60%.
+As a result of retraining the above five detectors on the sub-dataset Text2Video-zero, their performance on seen video generation models greatly improved, but they still performed poorly for unseen generation models, with ACC and AUC scores still lower than 60%, and the retrained STCNN presents similar results.
 On the contrary, our DeCoF has excellent generalizability for various unseen generation models.
 Compared with previous video classification methods based on pretrained CLIP [24], our method has still made significant progress.
-Specifically, on the task of generated video detection, DeCoF outperforms previous detectors by 13.77% $\sim$ 31.38% in ACC and 8.44% $\sim$ 49.18% in AP.
+Specifically, on the task of generated video detection, DeCoF outperforms previous detectors by 13.77% $\sim$ 31.38% in ACC and 8.44% $\sim$ 49.18% in AUC.
 We also conducted cross-dataset evaluations.
@@
 The upper row represents the robustness of Gaussian blur, and the lower row represents the robustness of JPEG compression.
-We test the detector trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AP (%) report for robustness comparison.
+We test the detector trained on different sub-datasets (represented by different colors) on different sub-datasets and use the AUC (%) report for robustness comparison.
 Q2. How does our detector maintain robustness against noise and disturbances?
@@
 Figure 9 shows the performance of our model trained on different sub-datasets in the face of the above two perturbations.
-Although experiencing some performance inconsistencies due to differences between training datasets, our method remains effective under all degrees of arbitrary perturbation, with the AP higher than 80%.
+Although experiencing some performance inconsistencies due to differences between training datasets, our method remains effective under all degrees of arbitrary perturbation, with the AUC higher than 80%.
 Q3. How does our detector perform on commercial generation models?
@@
 Gen-2 Pika Sora Veo
-ACC AP ACC AP ACC AP ACC AP
+ACC AUC ACC AUC ACC AUC ACC AUC
 Text2Video-zero 99.49 100.00 99.49 99.98 96.87 99.70 96.43 100.00
@@
 To further evaluate the effectiveness of our detection method, we constructed four additional test datasets using well-known commercial proprietary T2V models, Pika [32], Gen-2 [15], Sora [41] and Veo [45].
-As shown in Table 3, the detector achieved at least 94.84% ACC and 99.52% AP on Pika and at least 98.46% ACC and 99.96% AP on Gen-2.
+As shown in Table 3, the detector achieved at least 94.84% ACC and 99.52% AUC on Pika and at least 98.46% ACC and 99.96% AUC on Gen-2.
 Due to regional restrictions, we did not qualify for testing Sora and Veo directly, but we constructed test sets for 48 public videos released by Sora and 14 public videos released by Veo, respectively.
-In this test, we achieved no less than 89.58% ACC and no less than 97.92% AP on Sora, no less than 67.86% ACC and no less than 94.90% AP on Veo, demonstrating the strong generalizability of our model DeCoF for unseen commercial T2V models.
+In this test, we achieved no less than 89.58% ACC and no less than 97.92% AUC on Sora, no less than 67.86% ACC and no less than 94.90% AUC on Veo, demonstrating the strong generalizability of our model DeCoF for unseen commercial T2V models.
 7 Conclusions and Future Work
@@
 We hope this study will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AIGC applications.
-Appendix A Appendix
-This supplementary material provides:
-  • Sec. B: A detailed description of dataset construction, including dataset structure diagrams, as well as specific examples of prompts and videos.
-  • Sec. C : Various parameter details for model training, hyperparameter settings, etc.
-  • Sec. D : The impact of selecting different mapping functions.
-  • Sec. E : The impact of video frame count.
-  • Sec. F : The impact of video frame sampling interval.
-  • Sec. G : More details on Sora.
-  • Sec. H : Spatial artifacts vs Temporal artifacts.
-  • Sec. I : Impact Statements.
-  • Sec. J : Limitation.
-  • Sec. K : More visual examples.
-Appendix B $GeneratedVideoForensics$ Dataset
-B.1 Dataset Composition
-The following section provides a more detailed description of how the $GeneratedVideoForensics$ was constructed and evaluated. The entire GVF dataset is generated in two steps:
-1) Selection of prompts and real videos: The prompts and real videos of the GVF dataset come from existing open-domain text-to-video datasets: MSVD  [6] and MSR-VTT  [50] datasets.
-When selecting samples, we followed the work of FETV  [25], Considering the collection of prompts from the above four aspects to simulate the sample distribution in real scenarios.
-More specifically, in the work of FETV, they classify each prompt from two aspects: the main content it describes and the attributes it controls during video generation.
-The spatial major content of the prompt can be divided into people, animals, plants, food, vehicles, buildings, artifacts, scenery, and illustrations.
-The corresponding spatial attribute controls are divided into quantity, color, and camera view.
-Likewise, the major contents and attribute controls of temporal are actions, kinetic motions, light change, and fluid motions, respectively, motion direction, speed, and event sequence.
-2) Selection of generation models: Although there are many high-quality generation models on the market, we do not want to choose the best.
-On the contrary, we have selected four generation models with their advantages, aiming to simulate the crises we are encountering in real life: a large number of generation models appear, how do we deal with the problems caused by unseen generation models?
-In the end, the four most representative generation models were selected as our generation models: Text2Video-zero [22], ModelScopeT2V [46], ZeroScope[42], Show1 [51].
-At the same time, we also selected the three most popular models as the secret test dataset: Pika [32], Gen-2 [15], Sora[41] and Veo [45].
-For Sora and Veo, we provide detailed information in the section G
-If divided by content, the $GenerateVideoForensics$ data set $D$ consists of 964 triples, which can be expressed as $D=\{(R,P,F)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent real videos and corresponding text prompts, $F$ represents multiple fake videos generated by different generation models based on text prompts $P$.
-If divided by the generation model, the GenerateVideoForensics data set can be divided into four sub-datasets: Text2Video-zero, ModelScopeT2V, ZeroScope, and Show-1.
-Each sub-dataset consists of 964 new triples, which can be expressed as $D=\{(R,P,f)^{(i)}\}_{i=1}^{964}$, $R$ and $P$ represent the real video and the corresponding text prompt, and $f$ represents the fake video generated by the specific generation model represented by the name of the current sub-dataset based on the text prompt $P$.
-B.2 Dataset Evaluation
-As part of this research, we present a comprehensive analysis of the GVF dataset, which includes an analysis of the attribute statistics of prompts, as well as an assessment of the different generation models in terms of the spatial content, spatial attributes, temporal content, and temporal attributes according to the FETV [25].
-Table 4: Some examples of prompts.
-Each prompt is composed of a random combination of four attributes: main content of time, main attributes of time, main content of space, and main attributes of space.
-Here we list some examples of prompts, red for content and blue for attributes.
-Main content Attribute
-control
-Example
-Spatial People Quantity four friends are driving in the car
-Animals Color a white bird chirps and plays on a handmade staircase
-Plants Camera View a beautiful view of greenery and river from a top view
-Food Color in cup full of chocolate then we add yellow cream to it so yummy
-Temporal Actions Event Order a red-headed woman is shown walking and then sitting
-Actions Speed a man fastly cross the road by run and get in to the car
-Kinetic Motions Motion Direction the big black tractor is moving away from it s carrier
-Fluid Motions Speed a bucket of slime is quickly poured on top of a woman’s head
-Attribute statistics of prompts.
-From the four aspects mentioned before, the main content of space and time and the main attributes of space and time, we provide attribute statistics of prompts of the data set.
-As the figure 10 shows, the various attributes of prompts are different because one prompt may have multiple attributes at the same time.
-At the same time, different attributes have different difficulties in generating videos, and the harm they cause in a trust crisis is also different.
-Here, we also provide some specific examples of prompts, As shown in the table 4.
-Figure 10: Data distribution over categories under the “major content” (upper) and “attribute control” (lower) aspects.
-Evaluation of different generation models.
-Here, we present our assessment of four generation models according to the FETV standard [25].
-We assess the generated videos based on four perspectives: (i) Static quality focuses on the visual quality of single video frames.
-(ii) Temporal quality focuses on the temporal coherence of video frames.
-(iii) Overall alignment measures the overall alignment between a video and the given text prompt.
-(iv) Fine-grained alignment measures the video-text alignment regarding specific attributes.
-As shown in figure 11, Regarding spatial categories, the performance quality of these four data sets varies, and the performance of the same data set differs significantly between the static quality and temporal quality levels.
-The videos that include people, animals, and vehicles have poor spatial quality, but the videos that include plants, scenery, and natural items have greater spatial quality.
-The reason behind this is that the videos in the first three categories contain numerous high-frequency intricacies, such as face and finger, which are more difficult to comprehend compared to the low-frequency information found in the final two categories, such as water and clouds.
-In terms of static quality, Text2Video-zero performs well but performs poorly in terms of temporal quality.
-Text2Video-zero inherited the ability to generate images from Stable Diffusion, but it lacks temporal coherence in its videos due to the lack of video training.
-Show-1 is second only to Text2Video-zero in terms of static quality, and its temporal quality is also superior.
-In comparison to ModelScopeT2V, ZeroScope is inferior to ModelScope in temporal quality but performs worse in static quality.
-The reason for this may be that ZeroScope has been primarily designed for videos with a 16:9 aspect ratio and without watermarks, rather than focusing on improving visual quality and aligning video text.
-It is important to note that while the four models have better control over color and camera view, they do not perform as well when it comes to accurately controlling attributes such as quantity, motion direction, and order of the events.
-In terms of speed and motion direction attribute control, ModelScope and ZeroScope are particularly promising, whereas Text2Video-zero demonstrates poor performance.
-On different attributes of the four models, fine alignment and global alignment perform similarly, although there are some exceptions.
-This is due to the fact that other properties may have an effect on the overall alignment.
-Therefore, it is more accurate to assess controllability using fine alignment when examining a specific attribute, whereas global alignment is a reasonable approximation.
-Figure 11: The manual evaluation provided results across major static and temporal contents and attributes to control video-text alignment.
-Appendix C Parameter Description
-In this section, we provide a detailed description of various parameters of the experiment, including the hyperparameters of the model and the details of its training.
-Training details.
-During the training process, we only train on one sub-dataset of the GVF dataset and test on the other sub-datasets.
-The numbers of the training set, validation set, and test set are 771, 96, and 97, respectively.
-We freeze the parameters of CLIP:ViT [8], making it a pure mapping function, and only train our model using the SGD optimizer with an initial learning rate of 0.001 and a momentum coefficient of 0.9.
-All experiments were completed on two RTX 3090 graphics cards.
-Data pre-processing.
-All the following experiments are performed on our GVF dataset.
-In order to better detect the videos generated by the unseen generate model, we only train the model on the sub-dataset text2video-zero and test on the remaining three sub-datasets Show1 [51], ModelScope [46] and ZeroScope [42] Since the resolution and length of the videos generated by the four generation models are inconsistent, We take eight frames of each video for training and testing and center-crop each frame according to its shortest side, then resize it to 224×224.
-During the training process, we first resized the image to 256×256 according to bilinear interpolation and then randomly cropped with the size of 224 × 224.
-Some data augmentation, such as Jpeg compression, random horizontal flipping, and Gaussian blur, are also used.
-Model hyperparameters.
-Our model consists of two layers of transformer and an MLP head is an adaptation of the sequence to sequence trans- former, where the dimension and feature dimension of MLP is 768, path_size is 32, the dropout parameter is 0.1, and the number of attention heads is 4.
-Appendix D Different Mapping Functions
-In this section, we examine the impact of different mapping functions.
-The mapping function should have two properties.
-One is that the distance between features in the feature space is inversely proportional to the similarity between images, and the other is that it is insensitive to spatial artifacts.
-The second requirement is easy to implement as long as the mapping function is not used to train the fake content detection task.
-For the first requirement, we explored several different mapping functions, as shown in the table 5.
-First, we discussed the two different tasks of ImageNet and CLIP.
-Compared to ImageNet’s image classification task, CLIP’s image encoder trained on text image task alignment is a better mapping function.
-We considered two reasons: one is the size of the training data, and the other is due to the difference in the task itself; We do not want the model to classify images into specific items or biological categories but rather cluster similar images.
-At the same time, we compared the impact of different model frameworks: Due to issues such as the number of model parameters, Resnet trained on the CLIP task is not suitable as a mapping function.
-Overall, the mapping function largely determines the effectiveness of our approach.
-Table 5: A comparison of the impact of different mapping functions.
-Mapping Function Pretrain Tasks Video generation models Total
-Avg.
-Text2Video-zero Show1 ModelScope ZeroScope
-ACC AP ACC AP ACC AP ACC AP ACC AP
-ViT-L/14 CLIP 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
-Resnet50 CLIP 96.92 99.82 52.60 80.66 57.22 80.19 50.00 67.65 64.19 82.08
-ViT-B/16 ImageNet 98.00 99.79 65.39 87.65 54.70 76.39 55.57 76.1 68.41 84.98
-Resnet50 ImageNet 99.5 99.96 54.2 81.9 51.43 67.97 49.93 58.65 63.76 77.12
-Appendix E The Impact of Video Frame Number
-Here we discuss the impact of the number of video frames used for detection.
-Table 6: A comparison of the impact of different frame numbers.
-Training dataset Frame Numbers Video generation models Total
-Avg.
-Text2Video-zero Show1 ModelScope ZeroScope
-ACC AP ACC AP ACC AP ACC AP ACC AP
-Text2Video-zero 8 99.50 100.00 71.50 95.83 78.25 96.13 76.50 96.13 81.44 97.02
-Show1 8 98.50 99.77 97.00 99.33 90.25 97.96 93.75 99.08 94.88 99.04
-ModelScope 8 98.50 99.94 80.25 95.68 97.00 99.53 93.75 98.93 92.38 98.52
-ZeroScope 8 97.00 99.81 75.50 96.52 75.75 96.42 98.50 99.91 86.69 98.17
-Text2Video-zero 16 99.50 99.99 60.70 95.13 65.55 93.55 63.27 95.38 72.25 96.01
-Show1 16 97.86 99.69 95.93 99.04 93.73 98.13 94.30 98.88 95.45 98.93
-ModelScope 16 100.00 100.00 81.25 95.41 98.36 99.55 95.36 99.19 93.74 98.53
-ZeroScope 16 97.43 99.83 80.75 96.10 82.18 96.26 96.93 99.72 89.32 97.97
-Text2Video-zero 24 99.50 100.00 57.57 95.14 55.7 94.10 60.18 96.32 68.23 96.39
-Show1 24 97.80 99.98 94.36 99.10 89.66 97.49 93.73 98.93 93.88 98.87
-ModelScope 24 100.00 100.00 74.68 95.57 97.00 99.32 90.73 99.30 90.60 98.54
-ZeroScope 24 98.43 99.96 83.32 97.00 85.02 96.56 97.50 99.78 91.06 98.32
-In our work, we selected 8 frames for generated video detection for two reasons: on the one hand, the current generation model still has this deficiency in generating long videos.
-Even the videos currently displayed by Sora can only generate about 300-600 frames, so we believe that 8 frames is an acceptable choice.
-On the other hand, in order to highlight the effectiveness of our method, we chose a challenging number of 8 frames, and a large number of experiments have also proven that our method indeed performs very well.
-We conducted experiments on the impact of frames 8, 16, and 24 on detection, as shown in the table 6.
-We found that the performance of 16 frames is generally better than that of 8 frames, except for the models trained on the Text2Video zero subdataset, which is consistent with our previous evaluation results of several generation models.
-Text2Video zero has the worst time continuity.
-In addition, the effect will become worse at 24 frames.
-One possible reason is that as the sequence increases, the temporal artifacts of the generated video become more unique, which weakens the model’s generalizability.
-Appendix F The Impact of Video Frame Sampling Interval.
-In this section, we discussed the impact of sampling interval on detection results. In previous experiments, we usually took the first 8 frames of the video for detection.
-Unlike previous experiments, this experiment was conducted on the latest generation model Sora, as Sora has made significant progress in generating frames compared to other generation models, making it easier for us to comprehensively discuss the impact of sampling strategies.
-As shown in the table 7, facing three different sampling strategies: 1) Take the first 8 frames; 2) Take 8 frames evenly from all frames; 3) Take 8 frames evenly from the first 300 frames; Our method did not show a significant decrease, which we attribute to two reasons: 1) The prompts we collected were comprehensive enough to cover the speed of video changes well.
-2) The duration of the test video is still around 10 seconds, so the impact of sampling interval is not very significant.
-For longer generated videos, the performance degradation caused by changes in sampling intervals will be more severe.
-Table 7: The impact of different sampling strategies
-Training dataset Sampling strategy Sora
-ACC AP
-Text2Video-zero First 8 frames 96.87 99.70
-Show1 First 8 frames 94.79 98.26
-ModelScope First 8 frames 94.79 98.83
-ZeroScope First 8 frames 89.58 97.92
-Text2Video-zero Take 8 frames evenly from all frames 96.87 99.48
-Show1 Take 8 frames evenly from all frames 94.79 98.61
-ModelScope Take 8 frames evenly from all frames 94.79 98.96
-ZeroScope Take 8 frames evenly from all frames 88.54 98.31
-Text2Video-zero Take 8 frames evenly from first 300 frames 97.91 99.65
-Show1 Take 8 frames evenly from first 300 frames 94.79 98.48
-ModelScope Take 8 frames evenly from first 300 frames 93.75 98.74
-ZeroScope Take 8 frames evenly from first 300 frames 90.62 98.22
-Appendix G More Details on Sora.
-G.1 Test data on Sora
-Due to regional restrictions, we were not eligible to test Sora.
-But we tested all of the 48 videos officially displayed by Sora.
-The test dataset consists of 48 official videos released by Sora and 48 real videos from the original test set, the test on Veo are similar.
-G.2 The detection performance of human testers on Sora
-We invited three testers to evaluate the authenticity of 96 videos (48 real and 48 Sora generated) from the previously mentioned Sora test set. The experimental results are shown in the table 8.
-Table 8: Compared to the test results of human testers
-Training dataset Sora
-TP+TN ACC
-Tester 1 57(48+9) 59.37
-Tester 2 66(47+19) 68.75
-Tester 3 60(48+12) 62.5
-Text2Video-zero 93 96.87
-Show1 91 94.79
-ModelScope 91 94.79
-ZeroScope 86 89.58
-The results of DeCoF in testing far exceed the level of human judgment.
-At the same time, we also need to point out that interviews with testers after testing revealed that their judgment on the videos generated by Sora was artificially high, as there were 4-5 flashy videos in the officially released videos that did not match the results that could appear in the real world, such as pirate ships in coffee cups.
-Meanwhile, according to the testers, their judgment is based on whether there is irregular movement (such as objects appearing out of thin air) or object threading (such as overlapping basketball and basket) and abnormal distortion (such as human fingers or limbs of moving animals) in the generated video.
-Appendix H Spatial artifacts vs Temporal artifacts.
-H.1 Detect generated videos from temporal artifacts
-Although our method attempts to detect generated videos from temporal artifacts, there is a question: is it possible that CLIP: ViT has differences in the features between the generated video and the real video frames when extracting image features, resulting in better detection performance?
-In other words, is our method really detect generated videos from temporal artifacts?
-To prove this, we repeated the previous probe experiment.
-As shown in the table 9, our model has also performed well in detecting test data without spatial artifacts.
-On the other hand, for the second set of test data in the previous probe experiment, our model has also successfully detected it.
-We believe that repeated stillness in multiple frames is also a type of temporal artifact to some extent.
-Table 9: The proof of our method detect generated videos from temporal artifact
-Method Video generation models
-Temporal Spatial
-ACC AP ACC AP
-I3D 51.17 51.61 86.67 98.61
-Slow 50.04 50.37 93.81 99.51
-X3D 50.04 54.31 78.39 92.84
-Mvit 50.48 55.29 72.16 92.59
-Ours 98.46 99.80 98.46 99.66
-H.2 Spatial artifacts vs Temporal artifacts.
-In this paper, we attempt to detect generated video from the perspective of temporal artifacts, but this does not mean that it is impossible to detect generated video from spatial artifacts.
-In fact, the difficulty of detecting spatial artifacts in generated videos lies in how to extract more general artifacts, as different generation models generate videos with different resolutions, color brightness, contrast, and so on.
-Appendix I Impact Statements
-Our research focuses on detecting generated videos.
-We developed the DeCoF model and shared our dataset to support detecting generated videos.
-These works are crucial to protect digital content and prevent the spread of disinformation.
-Nevertheless, these tools have the potential to be misused, resulting in a rivalry between video generation and detection technology.
-We want to advocate for the ethical use of technology and foster research in the creation of tools for verifying media authenticity.
-We are confident that this will result in protecting the public from the harms of disinformation, improving the clarity and authenticity of information distribution, and ensuring the protection of personal privacy for individuals.
-Appendix J Limitation
-In this paper, we tackle an overlooked challenge, i.e., generated video detection, by proposing a universal detector to deal with the trust crisis caused by various emerging text-to-video (T2V) models, especially unseen models.
-However, there are various methods of video forgery, such as Deepfake and malicious editing, in which our methods may experience significant performance degradation or inapplicability.
-Taking Deepfake detection as an example, our method does not add additional detection modules to the face.
-In addition, Deepfake videos are often locally tampered with, and the generated videos are generated as a whole.
-Our work focuses on detecting generated videos, which has a more diverse range of forgery scenarios and more flexible forgery techniques.
-Appendix K More Visual Examples
-Here, we provide more examples of generated videos. Figures 12 to Figures 15 are examples of real videos and generated videos corresponding to prompts.
-Figure 12: Real videos and videos generated by four generation models according to prompt: A person mixing ingredients in a mixing bowl.
-Figure 13: Real videos and videos generated by four generation models according to prompt: A kitten moves about.
-Figure 14: Real videos and videos generated by four generation models according to prompt: A young man is seated and playing a guitar.
-Figure 15: Real videos and videos generated by four generation models according to prompt: A boy rides a bike on a fence.